### v1

# Neural MJD: Neural Non-Stationary Merton Jump Diffusion for Time Series Prediction

Abstract
While deep learning methods have achieved strong performance in time series prediction, their black-box nature and inability to explicitly model underlying stochastic processes often limit their generalization to non-stationary data, especially in the presence of abrupt changes. In this work, we introduce Neural MJD, a neural network based non-stationary Merton jump diffusion (MJD) model. Our model explicitly formulates forecasting as a stochastic differential equation (SDE) simulation problem, combining a time-inhomogeneous Itô diffusion to capture non-stationary stochastic dynamics with a time-inhomogeneous compound Poisson process to model abrupt jumps. To enable tractable learning, we introduce a likelihood truncation mechanism that caps the number of jumps within small time intervals and provide a theoretical error bound for this approximation. Additionally, we propose an Euler-Maruyama with restart solver, which achieves a provably lower error bound in estimating expected states and reduced variance compared to the standard solver. Experiments on both synthetic and real-world datasets demonstrate that Neural MJD consistently outperforms state-of-the-art deep learning and statistical learning methods. Our code is available at https://github.com/YuanpeiG/neural-MJD.
1UBC; 2Vector Institute; 3University of Texas at Austin; 4Canada CIFAR AI Chair
1 Introduction
Real-world time series often exhibit a mix of continuous trends and abrupt changes (jumps) [1, 2]. For example, stock prices generally follow steady patterns driven by macroeconomic factors but can experience sudden jumps due to unexpected news or policy shifts [3, 4]. Similarly, retail revenue may rise seasonally but jump abruptly due to sales promotions or supply chain disruptions [5, 6]. These discontinuous changes pose significant challenges for temporal dynamics modeling.
Classical statistical models, e.g., Merton jump diffusion (MJD) [3] or more general Lévy processes [7], provide a principled approach for modeling such data with jumps. They are effective for small datasets with well-understood statistical properties [8, 9, 10]. However, their assumptions—such as independent and stationary increments—often fail in real-world non-stationary settings. Additionally, these models struggle to capture interdependencies across multiple time series, such as competition effects among colocated businesses [11] or spillover dynamics in stock markets driven by investor attention shifts [12, 13]. As a result, they are difficult to scale effectively to large datasets. In contrast, deep learning approaches have demonstrated strong empirical performance by effectively learning time-varying patterns from data [14, 15, 16, 17, 18, 19]. Despite their success, these models are often black-box in nature and lack explicit mathematical formulations to describe the underlying dynamics. This limits their interpretability and often results in poor generalization to non-stationary data with jumps.
To address these limitations, we propose Neural MJD, a neural parameterization of the non-stationary Merton jump diffusion model that combines the advantages of statistical and learning-based approaches. In particular, our contributions are as follows:
  • Our Neural MJD integrates a time-inhomogeneous Itô diffusion to capture non-stationary stochastic dynamics and a time-inhomogeneous compound Poisson process to model abrupt jumps. The parameters of the corresponding SDEs are predicted by a neural network conditioned on past data and contextual information.
  • To enable tractable learning, we present a likelihood truncation mechanism that caps the number of jumps within small time intervals and provide a theoretical error bound for this approximation. Additionally, we propose an Euler-Maruyama with restart solver for inference, which achieves a provably lower error bound in estimating expected states and reduced variance compared to the standard solver.
  • Extensive experiments on both synthetic and real-world datasets show that our model consistently outperforms deep learning and statistical baselines under both stochastic and deterministic evaluation protocols.
2 Related work
Many neural sequence models have been explored for time series prediction, e.g., long short-term memory (LSTM) [20], transformers [21], and state space models (SSMs) [22]. These models [23, 24, 25, 26, 27, 28, 29, 30, 31, 32] have shown success across domains, including industrial production [33], disease prevention [34, 35], and financial forecasting [36, 27]. To handle more contextual information, extensions incorporating spatial context have been proposed, e.g., spatio-temporal graph convolutional networks (STGCN) [37], diffusion convolutional recurrent neural networks (DCRNN) [38], and graph message passing networks (GMSDR) [39]. However, these models remain fundamentally deterministic and do not explicitly model stochastic temporal dynamics. Generative models, e.g., deep auto-regressive models [40] and diffusion/flow matching models [41, 42, 43], provide probabilistic modeling of the time series and generate diverse future scenarios [44, 45, 46, 47, 48, 49]. However, these models often face computational challenges, as either sampling or computing the likelihood can be expensive. Additionally, they do not explicitly model abrupt jumps, limiting their generalization ability to scenarios with discontinuities.
Another line of research integrates classical mathematical models, such as ordinary and stochastic differential equations (ODEs and SDEs), into deep learning frameworks [29, 50, 51, 52, 53, 54, 55]. In financial modeling, physics-informed neural networks (PINNs) [56] have been explored to incorporate hand-crafted Black-Scholes (BS) and MJD models as guidance to construct additional loss functions [57, 58]. However, these approaches differ from ours, as we directly parameterize the non-stationary MJD model using neural networks rather than imposing predefined model structures as constraints. Neural jump diffusion models have also been explored in the context of temporal point processes (TPPs), such as Hawkes and Poisson processes [59, 60]. However, these methods primarily focus on event-based modeling, where jumps are treated as discrete occurrences of events, thus requiring annotated jump labels during training. In contrast, our approach aims to predict time series values at any give time, irrespective of whether a jump occurs, without relying on labeled jump events. Moreover, since jump events are unknown in our setting, our likelihood computation is more challenging since it requires summing over all possible number of jumps.
Finally, various extensions of traditional MJD have been proposed in financial mathematics to handle non-stationary data [61, 62], such as the stochastic volatility jump (SVJ) model [63], affine jump models [64], and the Kou jump diffusion model [65]. However, these models rely on strong assumptions for analytical tractability, requiring manual design of parameter evolution and often being computationally expensive [66]. For example, the SVJ model combines Heston’s stochastic volatility with MJD under the assumption that volatility follows the Cox-Ingersoll-Ross (CIR) process, meaning it reverts to a fixed long-term mean. Despite this, it lacks a closed-form likelihood function. Moreover, variants with time-dependent parameters require calibrations on market data to obtain the functions of parameters [67]. In contrast, our model directly learns the parameters of the non-stationary MJD from data, which is similar to the classical SDE calibration for financial modeling, but provides better expressiveness and flexibility while still permitting closed-form likelihood evaluation.
3 Background
To better explain our method, we first introduce two prominent models in mathematical finance, the Black-Scholes model and the Merton jump diffusion model.
Black-Scholes (BS) model. The Black-Scholes model was developed by Fischer Black and Myron Scholes, assuming that asset prices follow a continuous stochastic process [68]. Specifically, the dynamics of asset price $S_{t}$ at time $t$ is described by the following SDE:
$\mathrm{d}S_{t}=S_{t}(\mu\mathrm{d}t+\sigma\mathrm{d}W_{t}),$ (1)
where $\mu$ is the drift rate, representing the expected return per unit of time, and $\sigma$ is the volatility, indicating the magnitude of fluctuations in the asset price. $W_{t}$ refers to a standard Wiener process.
Merton jump diffusion (MJD) model. To account for discontinuities in asset price dynamics, Robert C. Merton extended the BS model by introducing the MJD model [3]. This model incorporates an additional jump process that captures sudden and significant changes in asset prices, which cannot be explained by continuous stochastic processes alone.
The dynamics of the asset price $S_{t}$ in the MJD model are described by the following SDE:
$\mathrm{d}S_{t}=S_{t}((\mu-\lambda k)\mathrm{d}t+\sigma\mathrm{d}W_{t}+\mathrm{d}Q_{t}),$ (2)
where $Q_{t}$ follows a compound Poisson process and captures the jump part. Specifically, $Q_{t}=\sum^{N_{t}}_{i=1}(Y_{i}-1)$, where $Y_{i}$ is the price ratio caused by the $i$-th jump event occurring at the time $t_{i}$, i.e., $Y_{i}={S_{t_{i}}}/{S_{t_{i}^{-}}}$ and $N_{t}$ is the total number of jumps up to time $t$. $S_{t_{i}}$ and $S_{t_{i}^{-}}$ are the prices after and before the jump at time $t_{i}$, respectively. $Y_{i}$ captures the relative price jump size since $\mathrm{d}S_{t_{i}}/S_{t_{i}}=(S_{t_{i}}-S_{t_{i}^{-}})/S_{t_{i}^{-}}=Y_{i}-1$. The price ratio $Y_{i}$ follows a log-normal distribution, i.e., $\ln Y_{i}\sim{\mathcal{N}}(\nu,\gamma^{2})$, where $\nu$ and $\gamma^{2}$ are the mean and the variance respectively. $N_{t}$ denotes the number of jumps that occur up to time $t$ and follows a Poisson process with intensity $\lambda$, which is the expected number of jumps per unit time. To make the expected relative price change $\mathbb{E}[\mathrm{d}S_{t_{i}}/S_{t_{i}}]$ remain the same as in the BS model in Eq. 1, MJD introduces an additional adjustment in the drift term of the diffusion, i.e., $-\lambda k\mathrm{d}t$ in Eq. 2. In particular, we have,
$\mathbb{E}[\mathrm{d}Q_{t}]=\mathbb{E}[(Y_{N_{t}}-1)\mathrm{d}N_{t}]=\mathbb{E}[Y_{N_{t}}-1]\mathbb{E}[\mathrm{d}N_{t}],$
where we use the assumption of MJD that “how much it jumps" (captured by $Y_{N_{t}}$) and “when it jumps" (captured by $N_{t}$) are independent. For the log-normal distributed $Y_{N_{t}}$, we can compute the expected jump magnitude $\mathbb{E}[Y_{N_{t}}-1]=\exp{(\nu+{\gamma^{2}}/{2})}-1$. To simplify the notation, we define $k\coloneqq\mathbb{E}[Y_{N_{t}}-1]$. For the Poisson process $N_{t}$, we have $\mathbb{E}[\mathrm{d}N_{t}]=\lambda\mathrm{d}t$. Therefore, we have $\mathbb{E}[\mathrm{d}Q_{t}]=\lambda k\mathrm{d}t$, which justifies the adjustment term $-\lambda k\mathrm{d}t$ in Eq. 2.
The MJD model has an explicit solution for its asset price dynamics, given by:
$\ln\frac{S_{t}}{S_{0}}=\left(\mu-\lambda k-\frac{\sigma^{2}}{2}\right)t+\sigma W_{t}+\sum^{N_{t}}_{i=1}\ln Y_{i}.$ (3)
Based on this solution, the conditional probability of the log-return at time $t$, given the initial price $S_{0}$ and the number of jumps $N_{t}=n$, can be derived as:
$P\left(\ln S_{t}|S_{0},N_{t}=n\right)={\mathcal{N}}\left(a_{n},b_{n}^{2}\right),$ (4)
where $a_{n}=\ln S_{0}+\left(\mu-\lambda k-\frac{\sigma^{2}}{2}\right)t+n\nu$ and $b_{n}^{2}=\sigma^{2}t+\gamma^{2}n$. Therefore, we obtain the likelihood,
$P\left(\ln S_{t}|S_{0}\right)=\sum_{n=0}^{\infty}P\left(N_{t}=n\right)P\left(\ln S_{t}|S_{0},N_{t}=n\right)=\sum_{n=0}^{\infty}\frac{(\lambda t)^{n}}{n!}\frac{1}{\sqrt{2\pi b_{n}^{2}}}\exp\left(-\frac{(\ln S_{t}-a_{n})^{2}}{2b_{n}^{2}}\right).$ (5)
Here we use the fact that $P(N_{t}=n)$ follows a Poisson distribution. One can then perform the maximum likelihood estimation (MLE) to learn the parameters $\{\mu,\sigma,\lambda,\nu,\gamma\}$. Additionally, the conditional expectation has a closed-form,
$\mathbb{E}\left[S_{t}|S_{0}\right]$ $=S_{0}\exp(\mu t).$ (6)
The derivations of the above formulas are provided in App. A.
Figure 1: The overview of Neural MJD. Our model captures discontinuous jumps in time-series data and uncovers the underlying non-stationary SDEs from historical sequences and context information. Our method enables numerical simulations for future forecasting along time.
4 Methods
We consider the general time series prediction problem where, given the observed past data $\{S_{0},S_{-1},\dots,S_{-T_{p}}\}$, and optional contextual information (e.g., additional features) $C$, the goal is to predict the future values $\{S_{1},\dots,S_{T_{f}}\}$. Here $T_{p}$ and $T_{f}$ denote the past and future time horizons, respectively. While we use integer indices for denoting time, sampling time is not restricted to integers. Our model is built upon a diffusion process, which is inherently continuous in time and compatible with arbitrary sampling mechanism. An overview of our method is shown in Fig. 1.
4.1 Neural Non-Stationary Merton Jump Diffusion
In the vanilla MJD model, the increments from the Wiener and Compound Poisson processes are independent and stationary, i.e., $\sigma\mathrm{d}W_{t}\overset{\mathrm{iid}}{\sim}{\mathcal{N}}(0,\sigma^{2}\mathrm{d}t)$, $\mathrm{d}N_{t}\overset{\mathrm{iid}}{\sim}\mathrm{Pois}(\lambda\mathrm{d}t)$, and $\ln Y_{i}\overset{\mathrm{iid}}{\sim}{\mathcal{N}}(\nu,\gamma^{2})$. The stationary assumption may be too strong in the real-world applications. For example, the stock prices of certain companies, such as Nvidia, experienced significantly larger jumps in the past decade compared to the previous one. Specifically, we allow independent but non-stationary increments in MJD by introducing time-inhomogeneous parameters $\{\mu_{t},\sigma_{t},\lambda_{t},\gamma_{t},\nu_{t}\}_{0\leq t\leq T_{f}}$ in the original SDE. Thus, at any future time $t$, the modified SDE follows
$\mathrm{d}S_{t}$ $=S_{t}\left((\mu_{t}-\lambda_{t}k_{t})\mathrm{d}t+\sigma_{t}\mathrm{d}W_{t}+\int_{\mathbb{R}^{d}}(y-1)N(\mathrm{d}t,\mathrm{d}y)\right).$ (7)
Here $\lambda_{t}\in\mathbb{R}_{+}$ and $\sigma_{t},\mu_{t},k_{t}\in\mathbb{R}^{d}$, while $W_{t}$ denotes a $d$-dimensional standard Wiener process. With a slight abuse of notation, $\sigma_{t}\mathrm{d}W_{t}$ means element-wise product between two size-$d$ vectors. $N(\mathrm{d}t,\mathrm{d}y)$ is a Poisson random measure on $[0,T_{f}]\times\mathbb{R}^{d}$, which encodes both the timing and magnitude of jumps. Intuitively, $N(\mathrm{d}t,\mathrm{d}y)$ counts the number of jumps occurring within the infinitesimal time interval $[t,t+\mathrm{d}t]$ whose sizes fall within $[y,y+\mathrm{d}y]$. The jump component $S_{t}\int_{\mathbb{R}^{d}}(y-1)N(\mathrm{d}t,\mathrm{d}y)$ introduces abrupt discontinuities in the process, accounting for sudden shifts in data.
The statistical properties of the Poisson random measure are uniquely determined by its intensity measure $\lambda_{t}f_{Y}(t,y)\mathrm{d}y\mathrm{d}t$. In our model, the intensity measure controls time-inhomogeneous jump magnitudes and frequencies. Namely, jump times follow a Poisson process with time-dependent intensity $\lambda_{t}$ and jump magnitudes follow a time-dependent log-normal distribution, i.e., a jump $Y_{t}$ at time $t$ follows $\ln Y_{t}\sim\mathcal{N}(\nu_{t},\gamma_{t}^{2})$, where we denote the log-normal density of $Y_{t}$ as $f_{Y}(t,y)$. Similarly, we define $k_{t}\coloneqq\mathbb{E}[Y_{t}-1]=\exp(\nu_{t}+\frac{\gamma_{t}^{2}}{2})-1$ in the drift term. This makes Eq. 7 equivalent to using the compensated Poisson measure $\tilde{N}(\mathrm{d}t,\mathrm{d}y)\coloneqq N(\mathrm{d}t,\mathrm{d}y)-\lambda_{t}f_{Y}(t,y)\mathrm{d}y\mathrm{d}t$ to remove the expected contribution of jumps. Note Eq. 7 includes $k_{t},$ so that $\int_{\mathbb{R}^{d}}(y-1)\lambda_{t}f_{Y}(t,y)\mathrm{d}y\mathrm{d}t=\lambda_{t}\mathbb{E}[Y_{t}-1]\mathrm{d}t=\lambda_{t}k_{t}\mathrm{d}t.$ Namely, it can be rewritten as,
$\mathrm{d}S_{t}$ $=S_{t}\left(\mu_{t}\mathrm{d}t+\sigma_{t}\mathrm{d}W_{t}+\int_{\mathbb{R}^{d}}(y-1)\tilde{N}(\mathrm{d}t,\mathrm{d}y)\right).$
This preserves the martingale property of the process induced by $\tilde{N}(\mathrm{d}t,\mathrm{d}y)$, e.g., $\mathbb{E}[\mathrm{d}S_{t}/S_{t}]$ matches the drift term in the non-stationary Black–Scholes model without jumps.
More importantly, inspired by the amortized inference in VAEs [69], we use neural networks to predict these time-inhomogeneous parameters based on the historical data and the contextual information $C$,
$\mu_{t},\sigma_{t},\lambda_{t},\gamma_{t},\nu_{t}=f_{\theta}(S_{0},S_{-1},\dots,S_{-T_{p}},C,t),$ (8)
where $f$ is a neural network parameterized by $\theta$. To simplify the notation, we denote the set of all observed data as ${\mathcal{C}}\coloneqq\{S_{0},S_{-1},\dots,S_{-T_{p}},C\}$ from now on.
The stochastic process described by the SDE in Eq. 7 is formally an instance of an additive process [61, Ch. 14], characterized by independent but non-stationary increments. If $\sigma_{t},\mu_{t}\in L^{2}$, i.e., they are square-integrable functions, and $\max_{\tau}(\int_{\mathbb{R}^{d}}|y^{2}|\lambda_{\tau}f_{Y}(\tau,y)\mathrm{d}y)<\infty$, then our non-stationary MJD has a unique solution for every $S_{0}>0$ [61, Theorem 14.1]. As our prediction time horizon is a closed domain $t\in[0,T_{f}]$, these conditions are easily satisfied as long as the neural network $f_{\theta}$ does not produce unbounded values. At any future time $T$, the explicit solution of the SDE is given by,
$\ln\frac{S_{T}}{S_{0}}$ $=\int_{0}^{T}(\mu_{t}-\lambda_{t}k_{t}-\frac{\sigma_{t}^{2}}{2})\mathrm{d}t+\int_{0}^{T}\sigma_{t}\mathrm{d}W_{t}+\int_{0}^{T}\int_{\mathbb{R}^{d}}\ln y{N}(\mathrm{d}t,\mathrm{d}y).$ (9)
Next, we model the conditional probability of log-return $\ln S_{t}$ given initial price $S_{0}$ and past data ${\mathcal{C}}$,
$P\left(\ln S_{T}|S_{0},{\mathcal{C}}\right)=\sum_{n=0}^{\infty}\frac{\exp(-\int_{0}^{T}\lambda_{t}\mathrm{d}t)}{n!}\Phi_{n},$ (10)
where
$\Phi_{n}$ $=\int\cdots\int_{[0,T]}\Pi_{i=1}^{n}\lambda_{t_{i}}\phi(\ln S_{T};a_{n},b_{n}^{2})\,\mathrm{d}t_{1}\cdots\mathrm{d}t_{n},$
$a_{n}$ $=\ln S_{T}+\int_{0}^{T}\left(\mu_{t}-\lambda_{t}k_{t}-\frac{\sigma_{t}^{2}}{2}\right)\mathrm{d}t+\sum_{i=1}^{n}\nu_{t_{i}},\ b_{n}^{2}=\int_{0}^{T}\sigma_{t}^{2}\mathrm{d}t+\sum_{i=1}^{n}\gamma_{t_{i}}^{2}.$
Here $\phi(\ln S_{T};a_{n},b_{n}^{2})$ is the density of a normal distribution with mean $a_{n}$ and variance $b_{n}^{2}$. $t_{1:n}$ denote the timing of $n$ jumps. Further, we compute the conditional expectations as,
$\mathbb{E}\left[S_{T}|{\mathcal{C}}\right]=S_{0}\exp(\int_{0}^{T}\mu_{t}\mathrm{d}t).$ (11)
Please refer to App. B for derivations. Evaluating Eq. 9 and Eq. 10 is non-trivial due to time inhomogeneity and jumps, typically requiring Monte Carlo methods or partial integro-differential equation techniques for approximate solutions [61, Ch. 6, Ch. 12].
4.2 Tractable Learning Method
While Eq. 10 provides the exact likelihood function, evaluating it precisely is impractical due to 1) integrals with time-dependent parameters lacking closed-form solutions and 2) the infinite series over the number of jumps. To learn the model via the maximum likelihood principle, we propose a computationally tractable learning objective with parameter bootstrapping and series truncation.
First, given a finite number of future time steps $\{1,\dots,T_{f}\}$, we discretize the continuous time in SDEs to construct a piecewise non-stationary MJD. Our model predicts time-varying parameters $\{\mu_{\tau},\sigma_{\tau},\lambda_{\tau},\gamma_{\tau},\nu_{\tau}\}_{\tau=1}^{T_{f}}$. For any time $t\leq T_{f}$, we map it to an integer index via $\rho_{t}\coloneqq\lfloor t\rfloor+1$. Thus, the likelihood of the data at $t+\delta$ given the data at $t$, where $\rho_{t}-1\leq t<t+\delta<\rho_{t}$, is given by:
[…]
where $a_{n,\delta}=\ln S_{t}+(\mu_{\rho_{t}}-\lambda_{\rho_{t}}k_{\rho_{t}}-{\sigma_{\rho_{t}}^{2}}/{2})\delta+n\nu_{\rho_{t}}$ and $b_{n,\delta}^{2}=\sigma^{2}_{\rho_{t}}\delta+\gamma^{2}_{\rho_{t}}n$. This approach eliminates the need for numerical simulation to compute the integrals in Eq. 10 and has been widely adopted for jump process modeling [62, 70].
As for the conditional expectation, we have
$\mathbb{E}\left[S_{t}|{\mathcal{C}}\right]=S_{0}\exp(\sum_{j=1}^{\rho_{t}-1}\mu_{j}+(t-\rho_{t}+1)\mu_{\rho_{t}}).$ (12)
Derivation details are shown in App. B. Further, we jointly consider the likelihood of all future data:
$P(\ln S_{1},\cdots,\ln S_{T_{f}}|{\mathcal{C}})=\prod_{\tau=1}^{T_{f}}P(\ln S_{\tau}|\{\ln S_{j}\}_{j<\tau},{\mathcal{C}})=\prod_{\tau=1}^{T_{f}}P(\ln S_{\tau}|S_{\tau-1},{\mathcal{C}}),$
where we use the Markov property and the fact that $\ln(\cdot)$ is bijective.
Therefore, the MLE objective is given by:
[…]
Algorithm 1 Neural MJD Training
1:  repeat
2:    $({\mathcal{C}},S_{1:T_{f}})\sim\mathcal{D}_{\text{train}}$, with ${\mathcal{C}}=[S_{-T_{p}:0},C]$
3:    $\{\mu_{\tau},\sigma_{\tau},\lambda_{\tau},\nu_{\tau},\gamma_{\tau}\}_{\tau=1}^{T_{f}}\leftarrow f_{\theta}({\mathcal{C}})$
4:    $\hat{S}_{0}\leftarrow S_{0}$
5:    for $\tau=1,\cdots,T_{f}$ do
[…]
8:    Update $\theta$ via $-\nabla_{\theta}\sum_{\tau=1}^{T_{f}}\left(-\psi_{\tau}+\omega\|S_{\tau}-\hat{S}_{\tau}\|^{2}\right)$
9:  until converged
Algorithm 2 Euler-Maruyama with Restart Inference
0:  Solver step size $1/M$
1:  ${\mathcal{C}}\sim\mathcal{D}_{\text{test}}$, with ${\mathcal{C}}=[S_{-T_{p}:0},C]$
2:  $\{\mu_{\tau},\sigma_{\tau},\lambda_{\tau},\nu_{\tau},\gamma_{\tau}\}_{\tau=1}^{T_{f}}\leftarrow f_{\theta}({\mathcal{C}})$
3:  $t_{0}\leftarrow 0$, $N\leftarrow M\times T_{f}$
4:  for $i=1,\cdots,N$ do
5:    $t_{i}\leftarrow t_{i-1}+1/M,\rho_{t_{i}}\leftarrow\lfloor t_{i}\rfloor+1$
[…]
8:    $\zeta_{i}\leftarrow\kappa\nu_{\rho_{t_{i}}}+\sqrt{\kappa}\gamma_{\rho_{t_{i}}}z_{2}$
[…]
9:    if $(i-1)\text{ mod }M=0$ then
[…]
11:    else
12:      $\ln\bar{S}_{t_{i}}\leftarrow\ln\bar{S}_{t_{i-1}}+\alpha_{i}+\beta_{i}+\zeta_{i}$
13:  return $\{\bar{S}_{t_{i}}\}_{i=1}^{N}$
The training algorithm of our neural non-stationary MJD model is shown in Alg. 1. In computing the term $\ln P(\ln S_{\tau}|S_{\tau-1},{\mathcal{C}})$ of Eq. 13, instead of doing teacher forcing, we replace the ground truth $S_{\tau-1}$ with the conditional mean prediction $\mathbb{E}[S_{\tau-1}\mid{\mathcal{C}}]$ from Eq. 12. This design mitigates the discrepancy between training and inference while reducing error accumulation in sequential predictions, especially for non-stationary data. As shown in the ablation study in Sec. 5.3, this approach improves performance effectively. To further improve accuracy, we introduce an additional regularization term that encourages the conditional mean to remain close to the ground truth. Additionally, the for loop in Alg. 1 can be executed in parallel, as the conditional mean computation does not depend on sequential steps, significantly improving efficiency. Notably, our model imposes no restrictions on the neural network architecture, and the specific design details are provided in Sec. D.1.
Truncation error of likelihood function. Exact computation of $P(\ln S_{\tau}\mid S_{\tau-1},{\mathcal{C}})$ in Eq. 13 requires evaluating an infinite series, which is infeasible in practice. To address this, we truncate the series at order $\kappa\in\mathbb{N}_{+}$, i.e., limiting the maximum number of jumps between consecutive time steps. We establish the following theoretical result to characterize the decay rate of the truncation error:
Theorem 4.1.
Let the likelihood approximation error in Sec. 4.2, truncated to at most $\kappa$ jumps, be
$\Psi_{\kappa}(t,\delta)\coloneqq\sum_{n=\kappa+1}^{\infty}P\left(\Delta N=n\right)P\left(\ln S_{t+\delta}\mid S_{t},{\mathcal{C}},\Delta N=n\right).$
Then, $\Psi_{\kappa}(t,\delta)$ decays at least super-exponentially as $\kappa\to\infty$, with a convergence rate of $O(\kappa^{-\kappa})$.
The proof is provided in Sec. C.1. The truncation error is dominated by $\kappa$, with other time-dependent parameters absorbed into the big-$O$ notation. We set $\kappa$ to 5 to achieve better empirical performance.
4.3 Inference based on Euler Scheme
Once trained, our Neural MJD model enables simulations following Eq. 7 by computing the non-stationary SDE parameters with a single neural function evaluation (NFE) of $f_{\theta}$. Unlike models limited to point-wise predictions, Neural MJD supports continuous-time simulation across the entire future horizon. Although the training data consists of a finite set of time steps, $S_{1:T_{f}}$, the learned model can generate full trajectories from $t=0$ to $t=T_{f}$ at arbitrary resolutions.
Figure 2: Comparison of numerical simulations with and without restart strategy during inference.
The standard Euler-Maruyama (EM) method provides a general-purpose approach for simulating SDEs with simple implementation and proven convergence [71, 72]. However, MJD SDEs exhibit analytically derived variance that grows over time (see App. B), and the simulated trajectories produced using the vanilla EM, assuming sufficiently small error, reflect this growth as well. Notably, the resulting high-variance simulations can undermine the empirical reliability of future forecasts.
In our MJD model, it is possible to compute closed-form expressions for statistical properties such as the mean and variance at any point in time [61]; for instance, the analytical mean can be derived from Eq. 12. Building on this insight, we propose a hybrid analytic-simulation solver, the Euler-Maruyama with restart method, which periodically injects the exact analytical mean to improve accuracy and enhance stability, as shown in Alg. 2 and Fig. 2. Specifically, we discretize time using a uniform step size ${1}/{M}$ for simulation and set the restart points as the target times $\{1,\cdots,T_{f}\}$. The solver follows the standard EM method for Eq. 7 whenever a restart is unnecessary. Otherwise, it resets the state using the conditional expectation from Eq. 12.
Further, we prove that this restart strategy has a tighter weak-convergence error, particularly helpful for empirical forecasting tasks where the mean estimation is critical. Let $\epsilon_{t}\coloneqq|\mathbb{E}[g(\bar{S}_{t})]-\mathbb{E}[g(S_{t})]|$ be the standard weak convergence error [71], where $S_{t}$ is the ground truth state, $\bar{S}_{t}$ is the estimated one using certain sampling scheme and $g$ is a $K$-Lipschitz continuous function. We denote the weak convergence errors of our restarted solver and the standard EM solver by $\epsilon^{R}_{t}$ and $\epsilon^{E}_{t}$, respectively.
Proposition 4.2.
Let $1/M$ be the step size. Both standard EM and our solver exhibit a weak convergence rate of $O(1/M)$. Specifically, the vanilla EM has a weak error of $\epsilon_{t}^{E}\leq K\exp(Lt)/M$ for some constant $L>0$, while ours achieves a tighter weak error of $\epsilon_{t}^{R}\leq K\exp(L(t-\lfloor t\rfloor))/M$.
The proof and details are left to Sec. C.2. Our sampler is in the same spirit as the parareal simulation algorithms [73, 74, 75, 76]: it first obtains estimates at discrete steps and then runs fine-grained simulations for each interval. By resetting the state to the true conditional mean at the start of each interval, our sampler reduces mean estimation error and prevents error accumulation over time. Notably, the SDE simulation requires no additional NFEs and adds negligible computational overhead compared to neural-network inference, since it involves only simple arithmetic operations that can be executed efficiently on CPUs. For reference, we also present the standard EM solver in Sec. D.4.
5 Experiments
In this section, we extensively examine Neural MJD’s performance on synthetic and real-world time-series datasets, highlighting its applicability in business analytics and stock price prediction.
Table 1: Quantitative results on the synthetic dataset.
Mean Winner-takes-all Probabilistic
[…]
ARIMA 0.29 –0.15 N/A N/A
BS 0.25 0.02 0.20 0.12 0.22 0.08
MJD 0.21 0.08 0.18 0.15 0.20 0.09
XGBoost 0.18 0.17 N/A N/A
MLP 0.14 0.21 N/A N/A
NJ-ODE 0.15 0.20 N/A N/A
Neural BS 0.15 0.25 0.10 0.35 0.14 0.29
Neural MJD (ours) 0.09 0.32 0.07 0.39 0.09 0.34
Figure 3: Qualitative result on the synthetic dataset.
Baselines. We evaluate Neural MJD against a wide range of competitors, including statistical methods such as ARIMA [77], the BS model, and the MJD model. Additionally, we compare against learning-based approaches, including supervised learning models such as XGBoost [78], MLPs, and GCNs [79], as well as denoising diffusion models like DDPM [41], EDM [80], and flow matching (FM) [43]. We include recent neural ODE or SDE based learning methods such as NJ-ODE [81], NeuralCDE [29] and LatentSDE [55] for comparisons. We also design a baseline model, Neural BS, which shares the same architecture as Neural MJD but omits the jump component. For DDPM, EDM, FM, Neural BS, and Neural MJD, we share the same transformer-based backbone to ensure a fair comparison. Since some datasets contain graph-structured data as seen in the following section, we incorporate additional graph encoding steps based on Graphormer [82] to capture spatial features, which also justifies the inclusion of GCN as a baseline. Further details are provided in App. D.
Evaluation metrics. We employ Mean Absolute Error (MAE), Mean Squared Error (MSE), and the R-squared (R2) score as the primary evaluation metrics. To account for stochastic predictions, we run each stochastic models $10$ times and report results across three types of metrics: 1) Mean Metrics: used for deterministic models or to average the results of stochastic models; 2) Winner-Takes-All Metrics: we select the best prediction from ten stochastic samples to compute minMAE, minMSE, and maxR2; 3) Probabilistic Metrics: these metrics assess the likelihood of stochastic predictions and select the most probable outcome to calculate $p$-MAE, $p$-MSE, and $p$-R2. We mark N/A for inapplicable metrics for certain methods. Please refer to App. D for more details.
5.1 Synthetic Data
Data generation. We evaluate our algorithm on a scalar Merton jump diffusion model. The dataset consists of $N=10,000$ paths, generated using the standard EM scheme with 100 time steps. Using a sliding window with stride 1, we predict the next 10 frames from the past 10. The data is split into 60% training, 20% validation, and 20% testing. Refer to App. D for details.
Results. Tab. 1 reports quantitative results on the jump-driven synthetic dataset. Learning-based methods outperform traditional statistical models (ARIMA, BS, and MJD), and our Neural MJD tops all three evaluation protocols, surpassing Neural BS thanks to its explicit jump modeling objective. Qualitatively, as shown in Fig. 3, our Neural MJD generates larger, realistic jumps, while the baselines produce smoother but less accurate trajectories.
Table 2: Quantitative results on the SafeGraph&Advan business analytics dataset.
Metrics Mean Winner-takes-all Probabilistic
[…]
ARIMA 152.6 1.66e05 -0.183 N/A N/A
BS 135.5 1.05e05 0.102 112.8 9.01e04 0.159 121.5 9.87e04 0.138
MJD 131.8 9.98e04 0.127 109.6 8.48e04 0.169 117.6 9.02e04 0.144
XGBoost 124.0 9.76e04 0.303 N/A N/A
MLP 109.5 8.18e04 0.416 N/A N/A
GCN 95.2 7.12e04 0.432 N/A N/A
DDPM 68.5 4.75e04 0.501 58.9 4.48e04 0.529 N/A
EDM 57.6 4.35e04 0.525 49.4 3.76e04 0.556 N/A
FM 54.5 4.32e04 0.540 47.8 3.58e04 0.552 N/A
NeuralCDE 94.6 7.09e04 0.425 N/A N/A
LatentSDE 75.7 5.26e04 0.487 66.5 4.58e04 0.498 N/A
Neural BS 56.4 4.17e04 0.539 45.6 3.45e04 0.561 55.9 4.16e04 0.538
Neural MJD (ours) 54.1 4.18e04 0.549 42.3 3.19e04 0.565 53.0 4.10e04 0.550
5.2 Real-World Business and Financial Data
Business analytics dataset. The SafeGraph&Advan business analytics dataset combines proprietary data from Advan [83] and SafeGraph [84] to capture daily customer spending at points of interest (POIs) in Texas, USA. It includes time-series features (e.g., visits, spending) and static features (e.g., parking availability) for each POI, along with ego graphs linking each POI to its 10 nearest neighbors. Using a sliding window of 14 input days to predict the next 7, the dataset spans Jan.–Dec. 2023 for training, Jan. 2024 for validation, and Feb.–Apr. 2024 for testing.
Stock price dataset. The S&P 500 dataset [85] is a public dataset containing historical daily prices for 500 major US stocks. It comprises time-series data without additional contextual information. We construct a simple fully connected graph among all listed companies. Similarly to the business analytics dataset, we employ a sliding window approach with a stride of 1, using the past 14 days as input to predict the next 7 days. The dataset is divided into training (Jan.–Dec. 2016), validation (Jan. 2017), and testing (Feb.–Apr. 2017) sets. Refer to App. D for further details about the datasets.
Results. Tab. 2 reports results on the SafeGraph&Advan dataset covering POIs revenue prediction, which is measured in dollars. Denoising generative models (e.g., DDPM, EDM, FM) show strong performance, significantly outperforming simple supervised baselines like GCN. Neural MJD further improves upon the strong FM baseline, especially in winner-takes-all metrics, indicating better diversity and accuracy in generating plausible outcomes through simulated jumps. While denoising models support likelihood evaluation, their high computational cost—requiring hundreds of NFEs—makes them unsuitable for large datasets. In contrast, Neural MJD enables fast likelihood evaluation without such overhead, which enables the computation of probabilistic metrics.
Tab. 3 shows similar results on the S&P 500 dataset. FM again outperforms conventional baselines, including ODE based NJ-ODE and NeuralCDE, and Neural MJD achieves the best overall performance, effectively capturing volatility and discontinuities in stock time-series data.
Table 3: Quantitative results on the S&P 500 stock dataset.
Metrics Mean Winner-takes-all Probabilistic
[…]
ARIMA 62.1 3.67e04 -0.863 N/A N/A
BS 65.1 4.01e04 0.052 44.6 1.79e04 0.145 52.8 2.03e04 0.105
MJD 64.3 3.58e04 0.092 40.7 1.22e04 0.235 49.7 1.67e04 0.112
XGBoost 44.3 1.64e04 0.170 N/A N/A
MLP 44.4 1.57e04 0.205 N/A N/A
GCN 44.7 1.53e04 0.224 N/A N/A
NJ-ODE 46.8 1.69e04 0.208 N/A N/A
DDPM 42.2 1.88e04 0.235 36.8 8.42e03 0.470 N/A
EDM 37.1 1.68e04 0.249 27.6 5.01e03 0.542 N/A
FM 34.9 8.47e03 0.368 19.8 3.59e03 0.625 N/A
NeuralCDE 42.8 1.46e04 0.201 N/A N/A
LatentSDE 39.8 1.44e04 0.212 20.8 3.49e03 0.617 N/A
Neural BS 31.6 4.32e03 0.781 12.6 8.04e02 0.959 22.3 2.19e03 0.889
Neural MJD (ours) 16.8 1.22e03 0.938 5.2 1.61e02 0.992 15.0 1.02e03 0.948
5.3 Ablation Study
Table 4: Ablation study on the effect of teacher forcing (TF) and Euler–Maruyama (EM).
Mean Winner-takes-all Probabilistic
[…]
Ours 66.7 0.495 57.4 0.511 64.5 0.499
w. TF 101.5 0.325 85.6 0.331 99.8 0.324
w. EM 85.6 0.397 79.4 0.423 84.4 0.405
Table 5: Runtime comparison.
Model Train (ms) 1-run (ms) 10-run (ms)
MLP 65.2 52.2 N/A
GCN 271.3 250.7 N/A
FM 184.6 275.4 2696.3
Ours 183.5 166.8 179.2
We perform ablation studies to evaluate (i) the training algorithm described in Alg. 1 and (ii) the Euler-Maruyama with restart solver introduced in Sec. 4.3. For the ablations, we use 10% of the SafeGraph&Advan business analytics training set for training and evaluate on the full validation set. The results are presented in Tab. 5. Our training algorithm computes the MLE loss using the model predictions instead of ground truth, unlike teacher forcing. This improves training stability and reduces the generalization gap. Additionally, we empirically show that the vanilla EM solver results in higher variance and worse performance compared to our solver.
Additionally, we compare the runtime of our method against various baselines in Tab. 5. Thanks to the efficient numerical simulation-based forecasting framework that does not increase NFEs, our models are particularly well-suited for efficient multi-run stochastic predictions.
6 Conclusion
We introduced Neural MJD, a neural non-stationary Merton jump diffusion model for time series forecasting. By integrating a time-inhomogeneous Itô diffusion and a time-inhomogeneous compound Poisson process, our approach captures non-stationary time series with abrupt jumps. We further proposed a likelihood truncation mechanism and an improved solver for efficient training and inference respectively. Experiments demonstrate that Neural MJD outperforms state-of-the-art approaches. Future work includes extending to more challenging data types like videos.
7 Acknowledgements
This work was funded, in part, by the NSERC DG Grant (No. RGPIN-2022-04636), the Vector Institute for AI, Canada CIFAR AI Chair, a Google Gift Fund, and the CIFAR Pan-Canadian AI Strategy through a Catalyst award. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through the Digital Research Alliance of Canada alliance.can.ca, and companies sponsoring the Vector Institute www.vectorinstitute.ai/#partners, and Advanced Research Computing at the University of British Columbia. Additional hardware support was provided by John R. Evans Leaders Fund CFI grant. Y.L. and Y.G. are supported by the NSF grant IIS-2153468. Q.Y. is supported by UBC Four Year Doctoral Fellowships.
Appendix
[…]
l1
Appendix A Derivations of the Stationary Merton Jump Diffusion Model
In this section, we briefly review the mathematical derivations from classical textbooks to ensure the paper is self-contained. Our primary focus is on the case where the state variable $S$ is scalar, as is common in many studies. However, in Sec. 4, we extend our analysis to the more general $\mathbb{R}^{d}$ setting. Notably, in our framework, we do not account for correlations among higher-dimensional variables. For instance, the covariance matrix of the Brownian motion is assumed to be isotropic, meaning all components have the same variance. To maintain clarity and consistency with standard textbook conventions, we adopt scalar notations throughout this section for simplicity.
A.1 MJD and Lévy Process
Definition A.1.
Lévy process [61, Definition 3.1] A càdlàg (right-continuous with left limits) stochastic process $(X_{t})_{t\geq 0}$ on $(\Omega,\mathcal{F},\mathbb{P})$ with values in $\mathbb{R}^{d}$ such that $X_{0}=0$ is called a Lévy process if it possesses the following properties:
  1. 1.
    Independent increments: For every increasing sequence of times $t_{0},t_{1},\dots,t_{n}$, the random variables $X_{t_{0}},X_{t_{1}}-X_{t_{0}},\dots,X_{t_{n}}-X_{t_{n-1}}$ are independent.
  2. 2.
    Stationary increments: The law of $X_{t+h}-X_{t}$ does not depend on $t$.
  3. 3.
    Stochastic continuity: For all $\varepsilon>0$, $\lim_{h\to 0}\mathbb{P}(|X_{t+h}-X_{t}|\geq\varepsilon)=0.$
A Lévy process $(X_{t})_{t\geq 0}$ is a stochastic process that generalizes jump-diffusion dynamics, incorporating both continuous Brownian motion and discontinuous jumps. The Merton Jump Diffusion (MJD) model given by,
$\mathrm{d}S_{t}=S_{t}((\mu-\lambda k)\mathrm{d}t+\sigma\mathrm{d}W_{t}+\mathrm{d}Q_{t}),$ (14)
is a specific example of a Lévy process, as it comprises both a continuous diffusion component and a jump component. According to the Lévy–Itô decomposition [61, Proposition 3.7], any Lévy process can be expressed as the sum of a deterministic drift term, a Brownian motion component, and a pure jump process, which is represented as a stochastic integral with respect to a Poisson random measure.
A.2 Explicit Solution to MJD
To derive the solution to MJD in Eq. 2, based on [61, Proposition 8.14], we first apply Itô’s formula to the SDE:
$\mathrm{d}f(S_{t},t)$ $=\frac{\partial f(S_{t},t)}{\partial t}\mathrm{d}t+b_{t}\frac{\partial f(S_{t},t)}{\partial S_{t}}\mathrm{d}t+\frac{\omega_{t}^{2}}{2}\frac{\partial^{2}f(S_{t},t)}{\partial S_{t}^{2}}\mathrm{d}t+\omega_{t}\frac{\partial f(S_{t},t)}{\partial S_{t}}\mathrm{d}W_{t}$
$+[f(S_{t})-f(S_{t-})],$ (15)
where $b_{t}=(\mu-\lambda k)S_{t}$, $\omega_{t}=\sigma S_{t}$, and $S_{t-}$ represents the value of $S$ before the jump at time $t$.
By setting the function $f(S_{t},t)=\ln S_{t}$, the formula can be rearranged as:
$\mathrm{d}\ln S_{t}$ $=\frac{\partial\ln S_{t}}{\partial t}\mathrm{d}t+(\mu-\lambda k)S_{t}\frac{\partial\ln S_{t}}{\partial S_{t}}\mathrm{d}t+\frac{\sigma^{2}S_{t}^{2}}{2}\frac{\partial^{2}\ln S_{t}}{\partial S_{t}^{2}}\mathrm{d}t+\sigma S_{t}\frac{\partial\ln S_{t}}{\partial S_{t}}\mathrm{d}W_{t}$
$+[\ln(S_{t})-\ln(S_{t-})]$
$=(\mu-\lambda k)S_{t}\frac{1}{S_{t}}\mathrm{d}t+\frac{\sigma^{2}S_{t}^{2}}{2}(\frac{1}{-S_{t}^{2}})\mathrm{d}t+\sigma S_{t}(\frac{1}{S_{t}})\mathrm{d}W_{t}+[\ln(S_{t})-\ln(S_{t-})]$
$=(\mu-\lambda k)\mathrm{d}t-\frac{\sigma^{2}}{2}\mathrm{d}t+\sigma\mathrm{d}W_{t}+[\ln(S_{t})-\ln(S_{t-})]$ (16)
From the definition of the Compound Poisson process, we have that $S_{t}=Y_{i}S_{t^{-}}$, such that $\ln(S_{t})-\ln(S_{t^{-}})=\ln Y_{i}$. Here, $Y_{i}$ is the magnitude of the multiplicative jump. Therefore, integrating both sides of Eq. 16, we get the final explicit solution for MJD model:
$\ln S_{t}-\ln S_{0}=(\mu-\lambda k-\frac{1}{2}\sigma^{2})t+\sigma W_{t}+\sum^{N_{t}}_{i=1}\ln Y_{i}.$ (17)
We can reorganize the explicit solution as:
$S_{t}=S_{0}\exp\left(\left(\mu-\lambda k-\frac{\sigma^{2}}{2}\right)t+\sigma W_{t}+\sum^{N_{t}}_{i=1}\ln Y_{i}\right),$ (18)
since the drift term, diffusion term and jump term are independent, we can derive the mean of $S_{t}$ conditional on $S_{0}$:
$\mathbb{E}[S_{t}|S_{0}]=$ $S_{0}\mathbb{E}\left[\exp\left(\left(\mu-\lambda k-\frac{\sigma^{2}}{2}\right)t+\sigma W_{t}+\sum^{N_{t}}_{i=1}\ln Y_{i}\right)\right]$
$=$ $S_{0}\mathbb{E}\left[\exp\left(\left(\mu-\lambda k-\frac{\sigma^{2}}{2}\right)t\right)\right]\cdot\mathbb{E}\left[\exp\left(\sigma W_{t}\right)\right]\cdot\mathbb{E}\left[\exp\left(\sum^{N_{t}}_{i=1}\ln Y_{i}\right)\right]$
$=$ $S_{0}\exp\left(\left(\mu-\lambda k-\frac{\sigma^{2}}{2}\right)t\right)\cdot\exp\left(\frac{\sigma^{2}}{2}t\right)\cdot\exp\left(\lambda kt\right)$
$=$ $S_{0}\exp\left(\mu t\right)$ (19)
A.3 Likelihood Function of MJD
For the log-likelihood derivation, given the conditional probability in Eq. 4, the log-likelihood of the MJD model can be expressed as:
$\log P\left(\ln S_{t}|S_{0}\right)=$ $\log\sum_{n=0}^{\infty}P\left(N_{t}=n\right)P\left(\ln S_{t}|S_{0},N_{t}=n\right)$
$=$ $\log\sum_{n=0}^{\infty}\exp\left({-\lambda t}\right)\frac{(\lambda t)^{n}}{n!}\frac{1}{\sqrt{2\pi b_{n}^{2}}}\exp\left(-\frac{(\ln S_{t}-a_{n})^{2}}{2b_{n}^{2}}\right)$
$=$ $\log\sum_{n=0}^{\infty}\exp\left(-\lambda t+\log\frac{\left(\lambda t\right)^{n}}{n!}+\log\frac{1}{\sqrt{2\pi b_{n}^{2}}}+\left(-\frac{\left(\ln S_{t}-a_{n}\right)^{2}}{2b_{n}^{2}}\right)\right)$
$=$ $\log\sum_{n=0}^{\infty}\exp\left({-\lambda t}+n\log{\left(\lambda t\right)}-\log{n!\sqrt{2\pi}}-\frac{\log{b_{n}^{2}}}{2}-\frac{\left(\ln S_{t}-a_{n}\right)^{2}}{2b_{n}^{2}}\right),$ (20)
where $a_{n}=\ln S_{0}+\left(\mu-\lambda k-\frac{\sigma^{2}}{2}\right)t+n\nu$ and $b_{n}^{2}=\sigma^{2}t+\gamma^{2}n$. In maximum likelihood estimation (MLE), the initial asset price $S_{0}$ is assumed to be constant (non-learnable) and can therefore be excluded from optimization. The objective of MLE is to estimate the parameter set $\Theta=\{\mu,\sigma,\lambda,\gamma,\nu\}$ by maximizing the likelihood of the observed data under the estimated parameters. For the MJD model, the MLE objective is to determine the optimal parameters $\hat{\Theta}$. By omitting constant terms and expanding $s_{n}$ and $a_{n}$, the final expression of the MLE objective can be simplified as:
$\hat{\Theta}$ $=\arg\max_{\Theta}\log P(\ln S_{t}|S_{0})$
$=\arg\max_{\Theta}\log\sum_{n=0}^{\infty}\exp\Bigl{(}{-\lambda t}+n\log{\left(\lambda t\right)}-\frac{\log\left({\sigma^{2}t+n\gamma^{2}}\right)}{2}$
$-\frac{(\ln S_{t}-\ln S_{0}-\left(\mu-\frac{\sigma^{2}}{2}-\lambda k\right)t-n\nu)^{2}}{2\left(\sigma^{2}t+n\gamma^{2}\right)}\Bigl{)}$ (21)
According to [86], the Fourier transform can be applied to the Merton Jump Diffusion log-return density function. The characteristic function is then given by:
$\phi_{c}(\omega)=$ $\int^{\infty}_{-\infty}\exp(i\omega x)P(x)\mathrm{d}x$
$=$ $\exp\left[\lambda t\left\{\exp\left(i\omega\nu+\frac{\gamma^{2}\omega^{2}}{2}\right)-1\right\}+i\omega\left(\left(\mu-\frac{\sigma^{2}}{2}-\lambda k\right)t\right)-\frac{\sigma^{2}\omega^{2}}{2}t\right],$ (22)
where $x=\ln\frac{S_{t}}{S_{0}}$.
With simplification $\phi_{c}(\omega)=\exp\left[tg(\omega)\right]$, we can find the characteristic exponent, namely, the cumulant generating function (CGF):
$g(\omega)=\lambda\left\{\exp\left(i\omega\nu+\frac{\gamma^{2}\omega^{2}}{2}\right)-1\right\}+i\omega\left(\mu-\frac{\sigma^{2}}{2}-\lambda k\right)-\frac{\sigma^{2}\omega^{2}}{2},$ (23)
where $k=\exp{(\nu+{\gamma^{2}}/{2})}-1$.
The series expansion of CFG is:
$g(\omega)=i\omega k_{1}-\frac{\omega^{2}k_{2}}{2}+\frac{\omega^{3}k_{3}}{6}...$ (24)
According to [61, Proposition 3.13], the cumulants of the Lévy distribution increase linearly with $t$. Therefore, the first cumulant $k_{1}$ is the mean of the standard MJD:
$k_{1}(t)=\mathbb{E}\left[\ln\frac{S_{t}}{S_{0}}\right]=(\mu-\lambda k-\sigma^{2}/2+\lambda\nu)t$ (25)
The second cumulant $k_{2}$ is variance of the standard MJD, which is:
$k_{2}(t)=\text{Var}\left[\ln\frac{S_{t}}{S_{0}}\right]=\left(\sigma^{2}+\lambda(\gamma^{2}+\nu^{2})\right)t$ (26)
The corresponding higher moments can also be calculated as:
$\text{Skewness}=k_{3}(t)=$ $\lambda(3\gamma^{2}\nu+\nu^{3})t$ (27)
Excess $\text{Excess Kurtosis}=k_{4}(t)=$ $\lambda(3\gamma^{4}+6\nu^{2}\gamma^{2}+\nu^{4})t$ (28)
Appendix B Derivations of the Non-stationary Merton Jump Diffusion Model
B.1 Non-stationary MJD and Additive Process
Definition B.1.
Additive process [61, Definition 14.1] A stochastic process $(X_{t})_{t\geq 0}$ on $\mathbb{R}^{d}$ is called an additive process if it is càdlàg, satisfies $X_{0}=0$, and has the following properties:
  1. 1.
    Independent increments: For every increasing sequence of times $t_{0},t_{1},\dots,t_{n}$, the random variables $X_{t_{0}},X_{t_{1}}-X_{t_{0}},\dots,X_{t_{n}}-X_{t_{n-1}}$ are independent.
  2. 2.
    Stochastic continuity: For all $\varepsilon>0$, $\lim_{h\to 0}\mathbb{P}(|X_{t+h}-X_{t}|\geq\varepsilon)=0.$
In the non-stationary MJD model, given by,
$\mathrm{d}S_{t}$ $=S_{t}((\mu_{t}-\lambda_{t}k_{t})\mathrm{d}t+\sigma_{t}\mathrm{d}W_{t}+\int_{\mathbb{R}^{d}}(y-1)N(\mathrm{d}t,\mathrm{d}y)),$ (29)
the parameters governing drift, volatility, and jump intensity evolve over time, resulting in non-stationary increments. This violates the key stationarity property required for Lévy processes, as discussed in App. A. Consequently, the non-stationary MJD no longer falls within the Lévy process framework. Instead, according to the definition above, a stochastic process with independent increments that follow a non-stationary distribution is classified as an additive process. Similar to the relationship between the stationary MJD and the Lévy process, the non-stationary MJD can be viewed as a specific instance of an additive process. Thus, we can apply corresponding mathematical tools for additive processes to study the non-stationary MJD.
B.2 Explicit Solution to Non-stationary MJD
To derive the explicit solution to the non-stationary MJD, according to [61, Proposition 8.19], we have the Itô formula for semi-martingales:
$f(t,X_{t})-f(0,X_{0})$ $=\int^{t}_{0}\frac{\partial f(s,X_{s})}{\partial s}\mathrm{d}s+\int^{t}_{0}\frac{\partial f(s,X_{s-})}{\partial x}\mathrm{d}X_{s}+\frac{1}{2}\int^{t}_{0}\frac{\partial^{2}f(s,X_{s-})}{\partial x^{2}}\mathrm{d}[X,X]^{c}_{s}$
$+\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[f(s,X_{s})-f(s,X_{s-})-\triangle X_{s}\frac{\partial f(s,X_{s-})}{\partial x}].$ (30)
According to [61, Remark 8.3], for a function independent of time (i.e., $f(t,X_{t})=f(X_{t})$), when we have finite number of jumps, we can rewrite the above equation as:
$f(X_{t})-f(X_{0})$ $=\int^{t}_{0}f^{\prime}(X_{s-})\mathrm{d}X_{s}^{c}+\frac{1}{2}\int^{t}_{0}f^{\prime\prime}(X_{s-})\mathrm{d}[X,X]^{c}_{s}+\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[f(X_{s})-f(X_{s-})],$
where $X_{s}^{c}$ is the continuous part of $X_{s}$, and $[X,X]^{c}_{s}$ is the continuous quadratic variation of $X$ over the interval $[0,s]$.
In our case, let $X_{t}=S_{t}$, and define $f(t,X_{t})=\ln S_{t}$, the corresponding derivatives are $\frac{\partial f(t,X_{t})}{\partial X_{t}}=\frac{\partial\ln S_{t}}{\partial S_{t}}=\frac{1}{S_{t}}$, and $\frac{\partial^{2}f(t,X_{t})}{\partial X_{t}^{2}}=\frac{1}{-S_{t}^{2}}$. The dynamics of non-stationary MJD is defined by:
$\mathrm{d}S_{t}$ $=S_{t}\left(\mu_{t}\mathrm{d}t-\lambda_{t}k_{t}\mathrm{d}t+\sigma_{t}\mathrm{d}W_{t}+\int_{\mathbb{R}^{d}}(y-1){N}(\mathrm{d}t,\mathrm{d}y)\right).$
The continuous part of the quadratic variation of $S_{t}$ is $\mathrm{d}[S,S]^{c}_{s}=S^{2}_{t}\sigma^{2}_{t}\mathrm{d}t$. A jump at time $s$ is modeled as a multiplicative change $S_{s}=yS_{s-}$. Thus, the jump contribution is $\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[f(X_{s})-f(X_{s-})]=\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[\ln(yS_{s-})-\ln(S_{s-})]=\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[\ln y]$. Since the jump process is driven by a Poisson random measure ${N}(\mathrm{d}t,\mathrm{d}y)$ on $[0,t]\times\mathbb{R}^{d}$, we can rewrite the sum over all jump times as an integral with respect to this measure. When there are finitely many jumps on $[0,t]$, we have $\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[\ln y]=\int_{0}^{t}\int_{\mathbb{R}^{d}}\ln y{N}(\mathrm{d}s,\mathrm{d}y)$.
Based on [61, Ch. 14], even when the parameters (drift, volatility, jump intensity, etc.) are time-dependent, the non-stationary MJD remains a semi-martingale. Therefore, we can simplify the equation as follows for time $t$:
$\ln S_{t}-\ln S_{0}$ $=\int^{t}_{0}\frac{1}{S_{s}}S_{s}\left(\mu_{s}-\lambda_{s}k_{s}\right)\mathrm{d}s-\frac{1}{2}\int^{s}_{0}\frac{1}{S_{s}^{2}}S_{s}^{2}\sigma_{s}^{2}\mathrm{d}s+\int_{0}^{t}\sigma_{s}\mathrm{d}W_{s}$
$+\int_{[0,t]\times\mathbb{R}^{d}}\ln yN(\mathrm{d}s,\mathrm{d}y)$ (31)
Therefore, the explicit solution is:
$\ln\frac{S_{t}}{S_{0}}$ $=\int_{0}^{t}(\mu_{s}-\lambda_{s}k_{s}-\frac{\sigma_{s}^{2}}{2})\mathrm{d}s+\int_{0}^{t}\sigma_{s}\mathrm{d}W_{s}+\int_{0}^{t}\int_{\mathbb{R}^{d}}\ln y{N}(\mathrm{d}s,\mathrm{d}y).$ (32)
The only assumption needed for the derivation is the finite variation condition: $\int_{0}^{t}\int_{\mathbb{R}}|y|N(\mathrm{d}s,\mathrm{d}y)<\infty$. Based on the explicit solution for $S_{t}$, we can easily compute the conditional expectations as,
$\mathbb{E}\left[\ln(S_{t}/S_{0})\right]=\int_{0}^{t}(\mu_{s}-\lambda_{s}k_{s}-\frac{\sigma_{s}^{2}}{2})\mathrm{d}s+\int_{0}^{t}\int_{\mathbb{R}^{d}}\ln y\lambda_{s}f_{Y}(s,y)\mathrm{d}y\mathrm{d}s.$ (33)
and
$\mathbb{E}\left[S_{t}|S_{0}\right]=S_{0}\exp(\int_{0}^{t}\mu_{s}\mathrm{d}s),$ (34)
The variance can also be calculated as,
$\text{Var}\left[\ln(S_{t}/S_{0})\right]=\int_{0}^{t}\sigma_{s}^{2}\mathrm{d}s+\int_{0}^{t}\int_{\mathbb{R}^{d}}(\ln y)^{2}\lambda_{s}f_{Y}(s,y)\mathrm{d}y\mathrm{d}s.$ (35)
Given the results for the general time-inhomogeneous system, one can directly substitute the coefficients into the discrete formulation implemented in Sec. 4.2 to obtain the corresponding results.
B.3 Likelihood Function of Non-stationary MJD
Let $X_{t}=\ln S_{t}/\ln S_{0},{t\geq 0}$ be the log-return of the asset price $S_{t}$. Under the non-stationary MJD settings, $X_{t}$ is an additive process, therefore by the general property of additive process [61, Ch 14], the law of $X_{t}$ is infinitely divisible and its characteristic function is given by the Lévy–Khintchine formula:
$\mathbb{E}[\exp(iu\cdot X_{t})]=\exp\psi_{t}(u),$
where
$\psi_{t}(u)=-\frac{1}{2}u\cdot A_{t}u+iu\cdot\Gamma_{t}+\int_{\mathbb{R}^{d}}\eta(\mathrm{d}y)\left(e^{iu\cdot x}-1-iu\cdot x\right),$ (36)
where we have the integrated volatility term $A_{t}=\int_{0}^{t}\sigma_{s}\mathrm{d}s$, the integrated drift term $\Gamma_{t}=\int_{0}^{t}(\mu_{s}-\lambda_{s}k_{s})\mathrm{d}s$, and the Lévy measure $\eta(\mathrm{d}y)=\lambda_{t}f_{Y}(t,y)$.
Since the jumps follow a time-inhomogeneous Poisson random measure and the process is additive, we can denote the integrated intensity of jumps by $\Lambda(t)=\int_{0}^{t}\lambda_{s}\mathrm{d}s$, then the number of jumps $N_{t}$ in the time range $[0,t]$ is a Poisson distribution with this integrated jump intensity $\Lambda(t)$. When conditioning on $N_{t}$, we will have:
$P(N_{t}=n)=\frac{\exp(-\Lambda(t))\Lambda(t)^{n}}{n!}$ (37)
We now derive the conditional density $P(\ln S_{t}|N_{t}=n,S_{0})$, and here we can start with the case of one jump. When there is exactly one jump in $[0,t]$, the jump time $s_{1}$ is random. Given jump time $s_{1}$, in a time-inhomogeneous setting, the instantaneous probability of a jump at time $s_{1}$ is proportional to $\lambda_{s_{1}}$. According to the dynamics of non-stationary MJD, the continuous part of the log-return leads to a normal distribution with mean being $a_{1}=\ln S_{0}+\int^{s_{1}}_{0}\left(\mu_{s}-\lambda_{s}k_{s}-\frac{\sigma_{s}^{2}}{2}\right)\mathrm{d}s+\nu_{s_{1}}$, and variance being $b^{2}_{1}=\int_{0}^{s_{1}}\sigma_{s}^{2}\mathrm{d}s+\gamma^{2}_{s_{1}}$. Thus, the conditional density of $\ln S_{t}$ given one jump at time $s_{1}$ is $\phi(\ln S_{t};a_{1},b_{1}^{2})$, where $\phi(\cdot;a_{1},b_{1}^{2})$ denotes the Gaussian density with mean $a_{1}$ and variance $b_{1}^{2}$.
Since the jump could have occurred at any time in $[0,t]$, we must integrate over the possible jump time $s_{1}$. Therefore, the conditional density given $N_{t}=1$ is:
$P(\ln S_{t}|N_{t}=1,S_{0})=\frac{1}{\Lambda(t)}\int_{0}^{t}\lambda_{s_{1}}\phi(\ln S_{t};a_{1},b_{1}^{2})\,\mathrm{d}s_{1},$ (38)
where $\frac{1}{\Lambda(t)}$ normalizes the density.
When generalizing to the case of $N^{t}=n$, the conditional density $P(\ln S_{t}|N_{t}=n,S_{0})$ is defined via an integration over the $n$ jump times, with the jump times denoted by $0\leq s_{1},...,s_{n}\leq t$.
Because the process is time-inhomogeneous, the probability density that a jump occurs at a specific time $s_{i}$ is given by the instantaneous rate $\lambda_{s_{i}}$, therefore for a given set of jump times, the joint density for the jumps is proportional to $\Pi_{i=1}^{n}\lambda_{s_{i}}$. The conditional density can be written as:
$P(\ln S_{t}|N_{t}=n,S_{0})=\frac{1}{\Lambda(t)^{n}}\int\cdots\int_{[0,t]}\Pi_{i=1}^{n}\lambda_{s_{i}}\phi(\ln S_{t};a_{n},b_{n}^{2})\,\mathrm{d}s_{1}\cdots\mathrm{d}s_{n}$ (39)
Here $\phi(\ln S_{t};a_{n},b_{n}^{2})$ is the density of a normal distribution with mean $a_{n}$ and variance $b_{n}^{2}$, which are defined by:
$a_{n}$ $=\ln S_{0}+\int_{0}^{t}\left(\mu_{s}-\lambda_{s}k_{s}-\frac{\sigma_{s}^{2}}{2}\right)\mathrm{d}s+\sum_{i=1}^{n}\nu_{s_{i}}$
$b_{n}^{2}$ $=\int_{0}^{t}\sigma_{s}^{2}\mathrm{d}t+\sum_{i=1}^{n}\gamma_{s_{i}}^{2}$
For convenience, we may write the mixture term as
$\Phi_{n}$ $=\int\cdots\int_{[0,t]}\Pi_{i=1}^{n}\lambda_{s_{i}}\phi(\ln S_{t};a_{n},b_{n}^{2})\,\mathrm{d}s_{1}\cdots\mathrm{d}s_{n}$ (40)
Therefore, for the time-varying SDEs, the conditional probability of $\ln S_{t}$ is given by,
$P\left(\ln S_{t}|S_{0}\right)=\sum_{n=0}^{\infty}\frac{\exp(-\Lambda(t))}{n!}\Phi_{n}.$ (41)
Appendix C Proofs of Theorem and Proposition
C.1 Proof of Theorem 4.1
See 4.1 Before diving into the proof, we first introduce two important lemmas.
Lemma C.1 (Theorem 2 in [2]).
Let $Y\sim\mathrm{Pois}(m)$ be a Poisson-distributed random variable with mean $m$. Its distribution function is defined as $P(Y\leq k)\coloneqq\exp(-m)\sum_{i=0}^{k}\frac{m^{i}}{i!},$ with integer support $k\in\{0,1,\dots,\infty\}$. For $k=0$ and $k=\infty$, one has $P(Y\leq 0)=\exp(-m),P(Y\leq\infty)=1.$ For every other $k\in\{1,2,3,\dots\}$, the following inequalities hold:
$\Phi\left(\mathrm{sign}(k-m)\sqrt{2H(m,k)}\right)<P(Y\leq k)<\Phi\left(\mathrm{sign}(k+1-m)\sqrt{2H(m,k+1)}\right),$
where $H(m,k)$ is the Kullback-Leibler (KL) divergence between two Poisson-distributed random variables with respective means $m$ and $k$:
$H(m,k)=D_{\text{KL}}\left(\mathrm{Pois}(m)\|\mathrm{Pois}(k)\right)=m-k+k\ln\left(\frac{k}{m}\right).$
And $\Phi(x)$ is the cumulative distribution function (CDF) of the standard normal distribution and $\mathrm{sign}(\cdot)$ is the signum function.
Lemma C.1 is particularly helpful in our proof below. We also acknowledge its foundation in an earlier work [87], which provides many insights and a profound amount of valuable knowledge on its own.
Lemma C.2 (Bounds on the Standard Normal CDF).
The following upper bound for $\Phi(\cdot)$ holds when $x<0$:
$\Phi(x)<\frac{\phi(x)}{|x|},$
where $\phi(x)=\frac{\exp(-x^{2}/2)}{\sqrt{2\pi}}$ is the probability density function of the standard normal distribution.
Proof.
By the Mills’ ratio inequality for the Gaussian distribution [88], we have $01-\Phi(x)<\frac{\phi(x)}{x},\forall x>$ 0 . Using the identity $\Phi(-x)=1-\Phi(x)$ for $x>0$, we immediately obtain: $\Phi(-x)<\frac{\phi(x)}{x},\forall x>0.$ For $x<0$, substituting $-x$ into the previous bound and noting that $\phi(-x)=\phi(x)$, we obtain $\Phi(x)<\frac{\phi(x)}{|x|},\forall x<0.$
Proof of Theorem 4.1.
Proof.
The original likelihood objective in Sec. 4.2 is as follows:
$P\left(\ln S_{t+\delta}|S_{t},{\mathcal{C}}\right)=\sum_{n=0}^{\infty}P\left(\Delta N=n\right)P\left(\ln S_{t+\delta}|S_{t},{\mathcal{C}},\Delta N=n\right)$ (42)
$=\sum_{n=0}^{\infty}\exp\left({-\lambda_{\rho_{t}}\delta}\right)\frac{(\lambda_{\rho_{t}}\delta)^{n}}{n!}\phi\left(\ln S_{t+\delta};a_{n,\delta},b^{2}_{n,\delta}\right)$
$=\sum_{n=0}^{\infty}\exp\left({-\lambda_{\rho_{t}}\delta}\right)\frac{(\lambda_{\rho_{t}}\delta)^{n}}{n!}\frac{1}{\sqrt{2\pi b_{n,\delta}^{2}}}\exp\left(-\frac{(\ln S_{t+\delta}-a_{n,\delta})^{2}}{2b_{n,\delta}^{2}}\right)$
where $\delta$ is a small time change so that $\rho_{t}-1\leq t<t+\delta<\rho_{t}$, $a_{n,\delta}=\ln S_{t}+(\mu_{\rho_{t}}-\lambda_{\rho_{t}}k_{\rho_{t}}-{\sigma_{\rho_{t}}^{2}}/{2})\delta+n\nu_{\rho_{t}}$ and $s_{n,\delta}^{2}=\sigma^{2}_{\rho_{t}}\delta+\gamma^{2}_{\rho_{t}}n$.
We define the truncation error with a threshold $\kappa$ as:
$\Psi_{\kappa}(t,\delta)\coloneqq\sum_{n=\kappa+1}^{\infty}P\left(\Delta N=n\right)P\left(\ln S_{t+\delta}\mid S_{t},{\mathcal{C}},\Delta N=n\right).$ (43)
The second term $P\left(\Delta N=n\right)P\left(\ln S_{t+\delta}\mid S_{t},{\mathcal{C}},\Delta N=n\right)$ is a Gaussian density function and upper bounded by $\frac{1}{\sqrt{2\pi b_{n,\delta}^{2}}}$, so the truncation error $\Psi_{\kappa}(t,\delta)$ is bounded by:
$\Psi_{\kappa}(t,\delta)$ $\leq\sum_{n=\kappa+1}^{\infty}\frac{1}{\sqrt{2\pi b_{n,\delta}^{2}}}P\left(\Delta N=n\right)$ (Gaussian density bound)
$\leq\frac{1}{\sqrt{2\pi b_{\kappa+1,\delta}^{2}}}\sum_{n=\kappa+1}^{\infty}P\left(\Delta N=n\right)$ ($b_{\kappa,\delta}$ increases as $\kappa$ goes up)
$=\frac{1}{\sqrt{2\pi b_{\kappa+1,\delta}^{2}}}(1-\sum_{n=0}^{\kappa}P\left(\Delta N=n\right))$ (property of Poisson CDF)
$<\frac{1}{\sqrt{2\pi b_{\kappa+1,\delta}^{2}}}\left(1-\Phi(\mathrm{sign}(\kappa-\lambda_{\rho_{t}}\delta)\sqrt{2D_{\text{KL}}(\mathrm{Pois}(\lambda_{\rho_{t}}\delta)\|\mathrm{Pois}(\kappa))})\right)$ (Lemma C.1)
$=\frac{1}{\sqrt{2\pi b_{\kappa+1,\delta}^{2}}}\Phi(\mathrm{sign}(\lambda_{t}\delta-\kappa)\sqrt{2D_{\text{KL}}(\mathrm{Pois}(\lambda_{\rho_{t}}\delta)\|\mathrm{Pois}(\kappa))})$ (Gaussian CDF)
As stated above, the KL divergence between two Poisson distributions follows
$D_{\text{KL}}\left(\mathrm{Pois}(a)\|\mathrm{Pois}(b)\right)=a-b+b\ln(\frac{b}{a})$
Therefore,
$\Psi_{\kappa}(t,\delta)$ $<\frac{1}{\sqrt{2\pi b_{\kappa+1,\delta}^{2}}}\Phi\left(\mathrm{sign}(\lambda_{\rho_{t}}\delta-\kappa)\sqrt{2D_{\text{KL}}(\mathrm{Pois}(\lambda_{\rho_{t}}\delta)\|\mathrm{Pois}(\kappa))}\right),$
$=\frac{1}{\sqrt{2\pi(\sigma^{2}_{\rho_{t}}\delta+\gamma^{2}_{\rho_{t}}(\kappa+1))}}\Phi\left(\mathrm{sign}(\lambda_{\rho_{t}}\delta-\kappa)\sqrt{2(\lambda_{\rho_{t}}\delta-\kappa+\kappa\ln(\frac{\kappa}{\lambda_{\rho_{t}}\delta}))}\right)$
$=\frac{1}{\sqrt{2\pi(\sigma^{2}_{\rho_{t}}\delta+\gamma^{2}_{\rho_{t}}(\kappa+1))}}\Phi\left(\mathrm{sign}(\frac{\lambda_{\rho_{t}}\delta}{\kappa}-1)\sqrt{2(\lambda_{\rho_{t}}\delta-\kappa-\kappa\ln(\frac{\lambda_{\rho_{t}}\delta}{\kappa}))}\right)$
Intuitively, the truncation error decreases to zero as $\kappa$ approaches infinity. Below, we analyze the convergence rate. When $\kappa$ is sufficiently large, the term $\mathrm{sign}\left(\frac{\lambda_{\rho_{t}}\delta}{\kappa}-1\right)$ is negative. Consequently, the upper bound becomes:
$\Psi_{\kappa}(t,\delta)$ $<\frac{1}{\sqrt{2\pi(\sigma^{2}_{\rho_{t}}\delta+\gamma^{2}_{\rho_{t}}(\kappa+1))}}\Phi\left(-\sqrt{2(\lambda_{\rho_{t}}\delta-\kappa-\kappa\ln(\frac{\lambda_{\rho_{t}}\delta}{\kappa}))}\right)$
$<\frac{1}{\sqrt{2\pi(\sigma^{2}_{\rho_{t}}\delta+\gamma^{2}_{\rho_{t}}(\kappa+1))}}\frac{\phi\left(-\sqrt{2(\lambda_{\rho_{t}}\delta-\kappa-\kappa\ln(\frac{\lambda_{\rho_{t}}\delta}{\kappa})}\right)}{\sqrt{2(\lambda_{\rho_{t}}\delta-\kappa-\kappa\ln(\frac{\lambda_{\rho_{t}}\delta}{\kappa}))}}$ (Lemma C.2)
$=\frac{\exp{(-\lambda_{\rho_{t}}\delta+\kappa+\kappa\ln(\frac{\lambda_{\rho_{t}}\delta}{\kappa}))}}{2\pi\sqrt{2(\sigma^{2}_{\rho_{t}}\delta+\gamma^{2}_{\rho_{t}}(\kappa+1))(\lambda_{\rho_{t}}\delta-\kappa-\kappa\ln(\frac{\lambda_{\rho_{t}}\delta}{\kappa}))}}$
As $\kappa\rightarrow\infty$, the numerator is dominated by $\exp(-\kappa\ln\kappa)$, which decays super-exponentially (faster than any polynomial or exponential decay). The denominator consists of two components:
  • The first term, $\sqrt{\sigma^{2}_{\rho_{t}}\delta+\gamma^{2}_{\rho_{t}}(\kappa+1)}$, scales asymptotically as $\gamma_{\rho_{t}}\sqrt{\kappa}$.
  • The second term, $\sqrt{\left(\lambda_{\rho_{t}}\delta-\kappa-\kappa\ln\left(\frac{\lambda_{\rho_{t}}\delta}{\kappa}\right)\right)}$, scales as $\sqrt{\kappa\ln\kappa}$.
Combining all terms, the upper bound scales as:
$\frac{1}{2\pi\sqrt{2}\gamma_{\rho_{t}}}\cdot\frac{\exp(-\kappa\ln\kappa)}{\kappa\sqrt{\ln\kappa}}\sim\frac{\kappa^{-\kappa}}{\kappa\sqrt{\ln\kappa}}.$
The term $\kappa^{-\kappa}$ decays super-exponentially, while the denominator grows algebraically (as $\kappa\sqrt{\ln\kappa}$). The rapid decay of $\kappa^{-\kappa}$ dominates the polynomial growth in the denominator. The overall convergence rate is super-exponentially fast, at the rate of $O(\exp(-\kappa\ln\kappa))$ or equivalently $O(\kappa^{-\kappa})$.
Since the upper bound of $\Psi_{\kappa}(t,\delta)$ decays at the rate of $O(\kappa^{-\kappa})$ and $\Psi_{\kappa}(t,\delta)$ is strictly positive, this implies that the original quantity $\Psi_{\kappa}(t,\delta)$ must decay at least as fast as the upper bound. This completes the proof.
C.2 Proof of Proposition 4.2
See 4.2
Proof.
Here, we prove that this restart strategy has a tighter weak-convergence error than the standard EM solver. Recall that we let $\epsilon_{t}\coloneqq|\mathbb{E}[g(\bar{S}_{t})]-\mathbb{E}[g(S_{t})]|$ be the standard weak convergence error [71], where $S_{t}$ is the ground truth state, $\bar{S}_{t}$ is the estimated one using certain sampling scheme and $g$ is a $K$-Lipschitz continuous function. We denote the weak convergence errors of our restarted solver and the standard EM solver by $\epsilon^{R}_{t}$ and $\epsilon^{E}_{t}$, respectively.
Step 1: Standard EM Results on Time-Homogeneous MJD SDEs
Early works on jump-diffusion SDE simulations explored the weak error bounds, which we summarize as follows. For time-homogeneous MJD SDEs, the error term $\epsilon_{t}^{E}$ of the standard EM method is dominated by $\epsilon_{t}^{E}\leq K\exp(Lt)/M$. This is supported by the following: (a) Theorem 2.2 in [72] establishes the $O(1/M)$ rate; (b) Sec. 4-5 of [72] and Theorem 2.1 of [89] shows that $\epsilon_{t}^{E}$ grows exponentially regarding time with a big-$O$ factor $O(e^{K_{p}(t)})$. In particular, the time-dependent term in the error bound $e^{K_{p}(t)}$ used in the proof of [72] is rooted in their Lemma 4.1, which can be proven in a more general setting in [89]; e.g., Eq. (2.16) in [89] discusses concrete forms of $K_{p}(t)$, which can be absorbed into $O(e^{Lt})$ for some constant $L>0$. Lastly, the $K$-Lipschitz condition of the function $g$ provides the coefficient $K$ in the bound. For a detailed proof—which is more involved and not central to the design and uniqueness of our algorithm—we refer the reader to [71, 90]. When combining the above existing results from the literature, we can derive the error bound of $\epsilon_{t}^{E}\leq K\exp(Lt)/M$.
Step 2: Standard EM Results on Time-Inhomogeneous MJD SDEs
Our paper considers time-inhomogeneous MJD SDEs, with parameters fixed within each interval $[\tau-1,\tau)$ ($\tau\in\mathbb{N}$, $\tau\geq 1$). This happens to align with the Euler-Peano scheme for general time-inhomogeneous SDEs approximation. As a specific case of time-varying Lévy processes, our MJD SDEs retain the same big-$O$ bounds as the time-homogeneous case. Namely, the standard EM solver has the same weak convergence error $\epsilon_{t}^{E}\leq K\exp(Lt)/M$, as in the time-homogeneous MJD SDEs. This can be justified by extending Section 5 of [72] that originally proves the EM’s weak convergence for time-homogeneous Lévy processes. Specifically, the core technique lies in the Lemma 4.1 of [72], which, based on [89], is applicable to both time-homogeneous and Euler-Peano-style inhomogeneous settings (see Remark 3.3.3 in [89]). Therefore, equivalent weak convergence bounds could be attained by extending Lemma 4.1 of [72] with proofs from [89] thanks to the Euler-Peano formulation.
Step 3: Our Restarted EM Solver Error Bound
We now discuss the error bound for the restarted EM solver, $\epsilon_{t}^{R}$. Thanks to explicit solutions for future states $\{S_{1},S_{2},\ldots,S_{T_{f}}\}$, we can analytically compute their mean $\mathbb{E}[S_{\tau}|\mathcal{C}]$, $\tau\geq 1$, based on Eq. 12, which greatly simplifies the analysis. Using the restart mechanism in line 10 of Alg. 2, we ensure that $\mathbb{E}[\bar{S}_{\tau}|\mathcal{C}]$ from our restarted EM solver closely approximates the true $\mathbb{E}[S_{\tau}|\mathcal{C}]$ at restarting times. $\epsilon_{t}^{R}$ is significantly reduced when restart happens (when $t$ is an integer in our context for simplicity), then it grows again at the same rate as the standard EM method until the next restart timestep. This explains the $O(e^{t-\lfloor t\rfloor})$ difference in the error bounds of $\epsilon_{t}^{R}$ and $\epsilon_{t}^{E}$, where $\lfloor t\rfloor$ is the last restart time. Note that we could make the restart timing more flexible to potentially achieve a tighter bound in terms of weak convergence. However, this may affect the diversity of the simulation results, as the fidelity of path stochasticity could be impacted. ∎
Appendix D Experiment Details
Figure 4: Neural MJD training pipeline. The symbol $\rho$ represents the MJD parameters $\{\mu_{\tau},\sigma_{\tau},\lambda_{\tau},\nu_{\tau},\gamma_{\tau}\}$ in our model.
D.1 Baseline, Model Architecture, and Experiment Settings
For the statistical BS and MJD baselines, we assume a stationary process and estimate the parameters using a numerical MLE objective based on past sequences. For the other deep learning baselines, including DDPM, EDM, FM, Neural BS, and Neural MJD, we implement our network using the standard Transformer architecture [21]. All baseline methods are based on the open-source code released by their authors, with minor modifications to adapt to our datasets. Note that the technical term diffusion in the context of SDE modeling (e.g., Merton jump diffusion) should not be conflated with diffusion-based generative models [41]. While both involve SDE-based representations of data, their problem formulations and learning objectives differ significantly.
We illustrate the training loss computation pipeline for Neural MJD in Fig. 4. Notably, the loss computation can be processed in parallel across the future time-step horizon, eliminating the need for recursive steps during training. We normalize the raw data into the range of $[0,1]$ for stability and use a regularization weight $\omega=1.0$ during training. All experiments were run on NVIDIA A40 and A100 GPUs (48 GB and 80 GB VRAM, respectively).
D.2 Datasets Details
For all datasets, we normalize the input data using statistics computed from the training set. For non-denoising models, normalization maps the data to the range $[0,1]$. In contrast, for denoising models (DDPM, EDM, FM), we scale the data to $[-1,1]$ to align with standard settings used in image generation. Importantly, normalization coefficients are derived solely from the training set statistics. Further details on this process are provided below.
Synthetic Data. We generate synthetic data using a scalar Merton Jump Diffusion model. The dataset consists of $N=10,000$ paths over the interval $[0,1]$, simulated using the Euler scheme with 100 time steps. To facilitate time-series forecasting, we employ a sliding window approach with a stride of 1, where the model predicts the next 10 frames based on the previous 10. The dataset is divided into 60% training, 20% validation, and 20% testing. For each simulation, model parameters are randomly sampled from uniform distributions: $\mu\sim U(0.1,0.5)$, $\sigma\sim U(0.1,0.5)$, $\lambda\sim U(3,10)$, $\nu\sim U(-0.1,0.1)$, and $\gamma\sim U(0.5,1.0)$. These parameter choices ensure the presence of jumps, capturing the stochastic nature of the process.
SafeGraph&Advan Business Analytics Data. The SafeGraph&Advan business analytics dataset is a proprietary dataset created by integrating data from Advan [83] and SafeGraph [84] to forecast daily customer spending at points of interest (POIs) across Texas, USA. Both datasets are licensed through Dewey Data Partners under their proprietary commercial terms, and we comply fully with the terms. For each POI, the dataset includes time-series data with dynamic features and static attributes. Additionally, ego graphs are constructed based on geodesic distances, where each POI serves as a central node connected to its 10 nearest neighbors. An visualization is shown in Fig. 5. Specifically, we use POI area, brand name, city name, top and subcategories (based on commercial behavior), and parking lot availability as static features. The dynamic features include spending data, visiting data, weekday, opening hours, and closing hours. These features are constructed for both ego and neighboring nodes. Based on the top category, we determine the maximum spending in the training data and use it to normalize the input data for both training and evaluation, ensuring a regularized numerical range. For training stability, we clip the minimum spending value to $0.01$ instead of 00 to enhance numerical stability for certain methods.
We adopt a sliding window approach with a stride of 1, using the past 14 days as input to predict spending for the next 7 days. The dataset spans multiple time periods: the training set covers January–December 2023, the validation set corresponds to January 2024, and the test set includes February–April 2024. This large-scale dataset consists of approximately 3.9 million sequences for training, 0.33 million for validation, and 0.96 million for testing.
Figure 5: Visualization of Ego Graph Dataset Construction in Austin, Texas
S&P 500 Stock Price Data. The S&P 500 dataset [85] is a publicly available dataset from Kaggle that provides historical daily stock prices for 500 of the largest publicly traded companies in the U.S (CC0 1.0 Universal license). It primarily consists of time-series data with date information and lacks additional contextual attributes. We include all listed companies and construct a simple fully connected graph among them. Therefore, for models capable of handling graph data, such as GCN, our implemented denoising models, and Neural MJD, we make predictions for all companies (represented as nodes) simultaneously. This differs from the ego-graph processing used in the SafeGraph&Advan dataset, where predictions are made only for the central node, while neighbor nodes serve purely as contextual information. To normalize the data, we determine the maximum stock price for each company in the training data, ensuring that input values fall within the $[0,1]$ range during training.
Following the approach used for the business analytics dataset, we apply a sliding window method with a stride of 1, using the past 14 days as input to predict stock prices for the next 7 days. The dataset is split into training (Jan.–Dec. 2016), validation (Jan. 2017), and testing (Feb.–Apr. 2017) sets. In total, it contains approximately 62K sequences for training, 5K for validation, and 15K for testing. To better distinguish the effects of different methods on the S&P 500 dataset, we use a adjusted R2 score $R^{2}=1-(1-R^{2}_{\text{reg}})\cdot\frac{n-1}{n-p-1}$, where $n$ is the sample size and we set the number of explanatory variables $p$ to be $(k-1)(n-1)/k$, where $k=70.0$.
D.3 Limitations
Our approach explicitly models discontinuities (jumps) in the time series. Consequently, if the underlying data lack such jump behaviors—i.e., if they are extremely smooth and exhibit no abrupt shifts—our jump component may be inaccurately estimated or effectively unused. In these scenarios, the model can underperform compared to simpler or purely continuous alternatives that do not rely on capturing sudden changes. For applications where jumps are absent or extremely rare, users should first verify the presence (or likelihood) of discontinuities in their dataset before adopting our framework. Additionally, one potential extension is to design an adaptive mechanism that can automatically deactivate or regularize the jump component when the data do not exhibit significant jump behavior, thereby reducing unnecessary complexity and improving general performance on smooth series.
D.4 Vanilla Euler Solver
Algorithm 3 Vanilla Euler-Maruyama Method
0:  Total solver steps $M$
1:  ${\bm{C}}\sim\mathcal{D}_{\text{test}}$, with ${\bm{C}}=[S_{-T_{p}:0},C]$
2:  $\{\mu_{\tau},\sigma_{\tau},\lambda_{\tau},\nu_{\tau},\gamma_{\tau}\}_{\tau=1}^{T_{f}}\leftarrow f_{\theta}({\bm{C}})$
[…]
4:  for $i=0,\cdots,M-1$ do
5:     $t_{i}\leftarrow i\Delta,t_{i+1}\leftarrow(i+1)\Delta,\rho_{t_{i}}\leftarrow\lfloor t_{i}\rfloor+1$
[…]
8:     $\zeta_{i}\leftarrow\kappa\nu_{\rho_{t_{i}}}+\sqrt{\kappa}\gamma_{\rho_{t_{i}}}z_{2}$
[…]
9:     $\ln\bar{S}_{t_{i+1}}\leftarrow\ln\bar{S}_{t_{i}}+\alpha_{i}+\beta_{i}+\zeta_{i}$
10:  return $\{\bar{S}_{t_{i}}\}_{i=1}^{M}$
We present the standard Euler–Maruyama solver above, which is used in the ablation study for comparison with our restarted Euler solver.
Appendix E Impact Statement
This paper introduces Neural MJD, a learning-based time series modeling framework that integrates principled jump-diffusion-based SDE techniques. Our approach effectively captures volatile dynamics, particularly sudden discontinuous jumps that govern temporal data, making it broadly applicable to business analytics, financial modeling, network analysis, and climate simulation. While highly useful for forecasting, we acknowledge potential ethical concerns, including fairness and unintended biases in data or applications. We emphasize responsible deployment and continuous evaluation to mitigate inequalities and risks.

### v2

@@
 Abstract
-While deep learning methods have achieved strong performance in time series prediction, their black-box nature and inability to explicitly model underlying stochastic processes often limit their generalization to non-stationary data, especially in the presence of abrupt changes.
+While deep learning methods have achieved strong performance in time series prediction, their black-box nature and inability to explicitly model underlying stochastic processes often limit their robustness handling non-stationary data, especially in the presence of abrupt changes.
 In this work, we introduce Neural MJD, a neural network based non-stationary Merton jump diffusion (MJD) model.
@@
 Experiments on both synthetic and real-world datasets demonstrate that Neural MJD consistently outperforms state-of-the-art deep learning and statistical learning methods.
-Our code is available at https://github.com/YuanpeiG/neural-MJD.
-1UBC; 2Vector Institute; 3University of Texas at Austin; 4Canada CIFAR AI Chair
+Our code is available at https://github.com/DSL-Lab/neural-MJD.
+1University of British Columbia; 2Vector Institute
+3University of Texas at Austin; 4Canada CIFAR AI Chair
 1 Introduction
@@
 This limits their interpretability and often results in poor generalization to non-stationary data with jumps.
+Notably, in previous deep learning time-series studies, “non-stationarity” typically refers to distributional shifts in the data over time.
+These works focus on mitigating such shifts using techniques like input-level normalization (e.g., DAIN [63], ST-norm [28], RevIN [42]), or domain adaptation (e.g., DDG-DA [50]).
+In contrast, our notion of “non-stationarity” centers on modeling a MJD process with parameters that evolve over time.
 To address these limitations, we propose Neural MJD, a neural parameterization of the non-stationary Merton jump diffusion model that combines the advantages of statistical and learning-based approaches.
@@
 $N_{t}$ denotes the number of jumps that occur up to time $t$ and follows a Poisson process with intensity $\lambda$, which is the expected number of jumps per unit time.
-To make the expected relative price change $\mathbb{E}[\mathrm{d}S_{t_{i}}/S_{t_{i}}]$ remain the same as in the BS model in Eq.
-1, MJD introduces an additional adjustment in the drift term of the diffusion, i.e., $-\lambda k\mathrm{d}t$ in Eq.
-2.
+To make the expected relative price change $\mathbb{E}[\mathrm{d}S_{t_{i}}/S_{t_{i}}]$ remain the same as in the BS model in Eq.˜1, MJD introduces an additional adjustment in the drift term of the diffusion, i.e., $\lambda k\mathrm{d}t$ in Eq.˜2.
 In particular, we have,
@@
 For the Poisson process $N_{t}$, we have $\mathbb{E}[\mathrm{d}N_{t}]=\lambda\mathrm{d}t$.
-Therefore, we have $\mathbb{E}[\mathrm{d}Q_{t}]=\lambda k\mathrm{d}t$, which justifies the adjustment term $-\lambda k\mathrm{d}t$ in Eq.
-2.
+Therefore, we have $\mathbb{E}[\mathrm{d}Q_{t}]=\lambda k\mathrm{d}t$, which justifies the adjustment term $\lambda k\mathrm{d}t$ in Eq.˜2.
 The MJD model has an explicit solution for its asset price dynamics, given by:
@@
 $\mathbb{E}\left[S_{t}|S_{0}\right]$ $=S_{0}\exp(\mu t).$ (6)
-The derivations of the above formulas are provided in App. A.
+The derivations of the above formulas are provided in App.˜A.
 Figure 1: The overview of Neural MJD.
@@
 Our model is built upon a diffusion process, which is inherently continuous in time and compatible with arbitrary sampling mechanism.
-An overview of our method is shown in Fig.
-1.
+An overview of our method is shown in Fig.˜1.
 4.1 Neural Non-Stationary Merton Jump Diffusion
@@
 With a slight abuse of notation, $\sigma_{t}\mathrm{d}W_{t}$ means element-wise product between two size-$d$ vectors.
-$N(\mathrm{d}t,\mathrm{d}y)$ is a Poisson random measure on $[0,T_{f}]\times\mathbb{R}^{d}$, which encodes both the timing and magnitude of jumps.
-Intuitively, $N(\mathrm{d}t,\mathrm{d}y)$ counts the number of jumps occurring within the infinitesimal time interval $[t,t+\mathrm{d}t]$ whose sizes fall within $[y,y+\mathrm{d}y]$.
+$N(\mathrm{d}t,\mathrm{d}y)$ is a Poisson random measure on $0,T_{f}]\times\mathbb{R}^{d}$, which encodes both the timing and magnitude of jumps.
+Intuitively, a Poisson random measure extends the idea of a Poisson process to random events distributed over both time and magnitude spaces, and $N(\mathrm{d}t,\mathrm{d}y)$ counts the number of jumps occurring within the infinitesimal time interval $t,t+\mathrm{d}t]$ whose sizes fall within $y,y+\mathrm{d}y]$.
 The jump component $S_{t}\int_{\mathbb{R}^{d}}(y-1)N(\mathrm{d}t,\mathrm{d}y)$ introduces abrupt discontinuities in the process, accounting for sudden shifts in data.
@@
 Similarly, we define $k_{t}\coloneqq\mathbb{E}[Y_{t}-1]=\exp(\nu_{t}+\frac{\gamma_{t}^{2}}{2})-1$ in the drift term.
-This makes Eq.
-7 equivalent to using the compensated Poisson measure $\tilde{N}(\mathrm{d}t,\mathrm{d}y)\coloneqq N(\mathrm{d}t,\mathrm{d}y)-\lambda_{t}f_{Y}(t,y)\mathrm{d}y\mathrm{d}t$ to remove the expected contribution of jumps.
-Note Eq.
-7 includes $k_{t},$ so that $\int_{\mathbb{R}^{d}}(y-1)\lambda_{t}f_{Y}(t,y)\mathrm{d}y\mathrm{d}t=\lambda_{t}\mathbb{E}[Y_{t}-1]\mathrm{d}t=\lambda_{t}k_{t}\mathrm{d}t.$ Namely, it can be rewritten as,
+This makes Eq.˜7 equivalent to using the compensated Poisson measure $\tilde{N}(\mathrm{d}t,\mathrm{d}y)\coloneqq N(\mathrm{d}t,\mathrm{d}y)-\lambda_{t}f_{Y}(t,y)\mathrm{d}y\mathrm{d}t$ to remove the expected contribution of jumps.
+Note Eq.˜7 includes $k_{t},$ so that $\int_{\mathbb{R}^{d}}(y-1)\lambda_{t}f_{Y}(t,y)\mathrm{d}y\mathrm{d}t=\lambda_{t}\mathbb{E}[Y_{t}-1]\mathrm{d}t=\lambda_{t}k_{t}\mathrm{d}t.$ Namely, it can be rewritten as,
 $\mathrm{d}S_{t}$ $=S_{t}\left(\mu_{t}\mathrm{d}t+\sigma_{t}\mathrm{d}W_{t}+\int_{\mathbb{R}^{d}}(y-1)\tilde{N}(\mathrm{d}t,\mathrm{d}y)\right).$
@@
 $\mu_{t},\sigma_{t},\lambda_{t},\gamma_{t},\nu_{t}=f_{\theta}(S_{0},S_{-1},\dots,S_{-T_{p}},C,t),$ (8)
-where $f$ is a neural network parameterized by $\theta$. To simplify the notation, we denote the set of all observed data as ${\mathcal{C}}\coloneqq\{S_{0},S_{-1},\dots,S_{-T_{p}},C\}$ from now on.
-The stochastic process described by the SDE in Eq.
-7 is formally an instance of an additive process [61, Ch.
+where $f$ is a neural network parameterized by $\theta$.
+To simplify the notation, we denote the set of all observed data as ${\mathcal{C}}\coloneqq\{S_{0},S_{-1},\dots,S_{-T_{p}},C\}$ from now on.
+Importantly, we only train a single network across all series and optimize the conditional log-likelihood as the training objective, which differs from standard statistical inference.
+At test time, the network produces context-dependent estimates $(\mu_{t},\sigma_{t},\lambda_{t},\nu_{t},\gamma_{t})$ across all future time in a single forward pass.
+The stochastic process described by the SDE in Eq.˜7 is formally an instance of an additive process [79, Ch.
 14], characterized by independent but non-stationary increments.
-If $\sigma_{t},\mu_{t}\in L^{2}$, i.e., they are square-integrable functions, and $\max_{\tau}(\int_{\mathbb{R}^{d}}|y^{2}|\lambda_{\tau}f_{Y}(\tau,y)\mathrm{d}y)<\infty$, then our non-stationary MJD has a unique solution for every $S_{0}>0$ [61, Theorem 14.1].
+If $\sigma_{t},\mu_{t}\in L^{2}$, i.e., they are square-integrable functions, and $\max_{\tau}(\int_{\mathbb{R}^{d}}|y^{2}|\lambda_{\tau}f_{Y}(\tau,y)\mathrm{d}y)<\infty$, then our non-stationary MJD has a unique solution for every $S_{0}>0$ [79, Theorem 14.1].
 As our prediction time horizon is a closed domain $t\in[0,T_{f}]$, these conditions are easily satisfied as long as the neural network $f_{\theta}$ does not produce unbounded values.
@@
 $\ln\frac{S_{T}}{S_{0}}$ $=\int_{0}^{T}(\mu_{t}-\lambda_{t}k_{t}-\frac{\sigma_{t}^{2}}{2})\mathrm{d}t+\int_{0}^{T}\sigma_{t}\mathrm{d}W_{t}+\int_{0}^{T}\int_{\mathbb{R}^{d}}\ln y{N}(\mathrm{d}t,\mathrm{d}y).$ (9)
-Next, we model the conditional probability of log-return $\ln S_{t}$ given initial price $S_{0}$ and past data ${\mathcal{C}}$,
-$P\left(\ln S_{T}|S_{0},{\mathcal{C}}\right)=\sum_{n=0}^{\infty}\frac{\exp(-\int_{0}^{T}\lambda_{t}\mathrm{d}t)}{n!}\Phi_{n},$ (10)
+Next, we model the conditional probability of log-return $\ln S_{t}$ given all the observed data $\mathcal{C}$},
+$P\left(\ln S_{T}|{\mathcal{C}}\right)=\sum_{n=0}^{\infty}\frac{\exp(-\int_{0}^{T}\lambda_{t}\mathrm{d}t)}{n!}\Phi_{n},$ (10)
 where
@@
 $\mathbb{E}\left[S_{T}|{\mathcal{C}}\right]=S_{0}\exp(\int_{0}^{T}\mu_{t}\mathrm{d}t).$ (11)
-Please refer to App.
-B for derivations.
-Evaluating Eq.
-9 and Eq.
-10 is non-trivial due to time inhomogeneity and jumps, typically requiring Monte Carlo methods or partial integro-differential equation techniques for approximate solutions [61, Ch.
+Please refer to App.˜B for derivations.
+Evaluating Eq.˜9 and Eq.˜10 is non-trivial due to time inhomogeneity and jumps, typically requiring Monte Carlo methods or partial integro-differential equation techniques for approximate solutions [79, Ch.
 6, Ch.
@@
 4.2 Tractable Learning Method
-While Eq.
-10 provides the exact likelihood function, evaluating it precisely is impractical due to 1) integrals with time-dependent parameters lacking closed-form solutions and 2) the infinite series over the number of jumps.
+While Eq.˜10 provides the exact likelihood function, evaluating it precisely is impractical due to 1) integrals with time-dependent parameters lacking closed-form solutions and 2) the infinite series over the number of jumps.
 To learn the model via the maximum likelihood principle, we propose a computationally tractable learning objective with parameter bootstrapping and series truncation.
@@
 Thus, the likelihood of the data at $t+\delta$ given the data at $t$, where $\rho_{t}-1\leq t<t+\delta<\rho_{t}$, is given by:
-[…]
+$\begin{aligned} P\left(\ln S_{t+\delta}|S_{t},{\mathcal{C}}\right)&=\sum_{n=0}^{\infty}P\left(\Delta N=n\right)P\left(\ln S_{t+\delta}|S_{t},{\mathcal{C}},\Delta N=n\right)\\ &=\sum_{n=0}^{\infty}\exp\left({-\lambda_{\rho_{t}}\delta}\right)\lambda_{\rho_{t}}^{n}\frac{\delta^{n}}{n!}\phi\left(\ln S_{t+\delta};a_{n,\delta},b^{2}_{n,\delta}\right),\end{aligned}$
 where $a_{n,\delta}=\ln S_{t}+(\mu_{\rho_{t}}-\lambda_{\rho_{t}}k_{\rho_{t}}-{\sigma_{\rho_{t}}^{2}}/{2})\delta+n\nu_{\rho_{t}}$ and $b_{n,\delta}^{2}=\sigma^{2}_{\rho_{t}}\delta+\gamma^{2}_{\rho_{t}}n$.
-This approach eliminates the need for numerical simulation to compute the integrals in Eq.
-10 and has been widely adopted for jump process modeling [62, 70].
+This approach eliminates the need for numerical simulation to compute the integrals in Eq.˜10 and has been widely adopted for jump process modeling [5, 26].
 As for the conditional expectation, we have
 $\mathbb{E}\left[S_{t}|{\mathcal{C}}\right]=S_{0}\exp(\sum_{j=1}^{\rho_{t}-1}\mu_{j}+(t-\rho_{t}+1)\mu_{\rho_{t}}).$ (12)
-Derivation details are shown in App. B. Further, we jointly consider the likelihood of all future data:
+Derivation details are shown in App.˜B. Further, we jointly consider the likelihood of all future data:
 $P(\ln S_{1},\cdots,\ln S_{T_{f}}|{\mathcal{C}})=\prod_{\tau=1}^{T_{f}}P(\ln S_{\tau}|\{\ln S_{j}\}_{j<\tau},{\mathcal{C}})=\prod_{\tau=1}^{T_{f}}P(\ln S_{\tau}|S_{\tau-1},{\mathcal{C}}),$
@@
 […]
-8:    Update $\theta$ via $-\nabla_{\theta}\sum_{\tau=1}^{T_{f}}\left(-\psi_{\tau}+\omega\|S_{\tau}-\hat{S}_{\tau}\|^{2}\right)$
+8:  Update $\theta$ via $\nabla_{\theta}\sum_{\tau=1}^{T_{f}}\left(-\psi_{\tau}+\omega\|S_{\tau}-\hat{S}_{\tau}\|^{2}\right)$
 9:  until converged
@@
 13:  return $\{\bar{S}_{t_{i}}\}_{i=1}^{N}$
-The training algorithm of our neural non-stationary MJD model is shown in Alg.
-1.
-In computing the term $\ln P(\ln S_{\tau}|S_{\tau-1},{\mathcal{C}})$ of Eq.
-13, instead of doing teacher forcing, we replace the ground truth $S_{\tau-1}$ with the conditional mean prediction $\mathbb{E}[S_{\tau-1}\mid{\mathcal{C}}]$ from Eq.
-12.
+The training algorithm of our neural non-stationary MJD model is shown in Alg.˜1.
+In computing the term $\ln P(\ln S_{\tau}|S_{\tau-1},{\mathcal{C}})$ of Eq.˜13, instead of doing teacher forcing, we replace the ground truth $S_{\tau-1}$ with the conditional mean prediction $\mathbb{E}[S_{\tau-1}\mid{\mathcal{C}}]$ from Eq.˜12.
 This design mitigates the discrepancy between training and inference while reducing error accumulation in sequential predictions, especially for non-stationary data.
-As shown in the ablation study in Sec.
-5.3, this approach improves performance effectively.
+As shown in the ablation study in Sec.˜5.3, this approach improves performance effectively.
 To further improve accuracy, we introduce an additional regularization term that encourages the conditional mean to remain close to the ground truth.
-Additionally, the for loop in Alg.
-1 can be executed in parallel, as the conditional mean computation does not depend on sequential steps, significantly improving efficiency.
-Notably, our model imposes no restrictions on the neural network architecture, and the specific design details are provided in Sec.
-D.1.
+Additionally, the for loop in Alg.˜1 can be executed in parallel, as the conditional mean computation does not depend on sequential steps, significantly improving efficiency.
+Notably, our model imposes no restrictions on the neural network architecture, and the specific design details are provided in Sec.˜D.1.
 Truncation error of likelihood function.
-Exact computation of $P(\ln S_{\tau}\mid S_{\tau-1},{\mathcal{C}})$ in Eq.
-13 requires evaluating an infinite series, which is infeasible in practice.
+Exact computation of $P(\ln S_{\tau}\mid S_{\tau-1},{\mathcal{C}})$ in Eq.˜13 requires evaluating an infinite series, which is infeasible in practice.
 To address this, we truncate the series at order $\kappa\in\mathbb{N}_{+}$, i.e., limiting the maximum number of jumps between consecutive time steps.
@@
 Theorem 4.1.
-Let the likelihood approximation error in Sec. 4.2, truncated to at most $\kappa$ jumps, be
+Let the likelihood approximation error in Sec.˜4.2, truncated to at most $\kappa$ jumps, be
 $\Psi_{\kappa}(t,\delta)\coloneqq\sum_{n=\kappa+1}^{\infty}P\left(\Delta N=n\right)P\left(\ln S_{t+\delta}\mid S_{t},{\mathcal{C}},\Delta N=n\right).$
 Then, $\Psi_{\kappa}(t,\delta)$ decays at least super-exponentially as $\kappa\to\infty$, with a convergence rate of $O(\kappa^{-\kappa})$.
-The proof is provided in Sec.
-C.1.
+The proof is provided in Sec.˜C.1.
 The truncation error is dominated by $\kappa$, with other time-dependent parameters absorbed into the big-$O$ notation.
@@
 4.3 Inference based on Euler Scheme
-Once trained, our Neural MJD model enables simulations following Eq.
-7 by computing the non-stationary SDE parameters with a single neural function evaluation (NFE) of $f_{\theta}$.
+Once trained, our Neural MJD model enables simulations following Eq.˜7 by computing the non-stationary SDE parameters with a single neural function evaluation (NFE) of $f_{\theta}$.
 Unlike models limited to point-wise predictions, Neural MJD supports continuous-time simulation across the entire future horizon.
@@
 The standard Euler-Maruyama (EM) method provides a general-purpose approach for simulating SDEs with simple implementation and proven convergence [45, 64].
-However, MJD SDEs exhibit analytically derived variance that grows over time (see App.
-B), and the simulated trajectories produced using the vanilla EM, assuming sufficiently small error, reflect this growth as well.
+However, MJD SDEs exhibit analytically derived variance that grows over time (see App.˜B), and the simulated trajectories produced using the vanilla EM, assuming sufficiently small error, reflect this growth as well.
 Notably, the resulting high-variance simulations can undermine the empirical reliability of future forecasts.
-In our MJD model, it is possible to compute closed-form expressions for statistical properties such as the mean and variance at any point in time [61]; for instance, the analytical mean can be derived from Eq.
-12.
-Building on this insight, we propose a hybrid analytic-simulation solver, the Euler-Maruyama with restart method, which periodically injects the exact analytical mean to improve accuracy and enhance stability, as shown in Alg.
-2 and Fig.
-2.
-Specifically, we discretize time using a uniform step size ${1}/{M}$ for simulation and set the restart points as the target times $\{1,\cdots,T_{f}\}$.
-The solver follows the standard EM method for Eq.
-7 whenever a restart is unnecessary.
-Otherwise, it resets the state using the conditional expectation from Eq.
-12.
+In our MJD model, it is possible to compute closed-form expressions for statistical properties such as the mean and variance at any point in time [79]; for instance, the analytical mean can be derived from Eq.˜12.
+Building on this insight, we propose a hybrid analytic-simulation solver, the Euler-Maruyama with restart method, which periodically injects the exact analytical mean to improve accuracy and enhance stability, as shown in Alg.˜2 and Fig.˜2.
+Specifically, we discretize time using a uniform step size 1/M{1}/{M} for simulation and set the restart points as the target times $\{1,\cdots,T_{f}\}$.
+The solver follows the standard EM method for Eq.˜7 whenever a restart is unnecessary.
+Otherwise, it resets the state using the conditional expectation from Eq.˜12.
 Further, we prove that this restart strategy has a tighter weak-convergence error, particularly helpful for empirical forecasting tasks where the mean estimation is critical.
@@
 Specifically, the vanilla EM has a weak error of $\epsilon_{t}^{E}\leq K\exp(Lt)/M$ for some constant $L>0$, while ours achieves a tighter weak error of $\epsilon_{t}^{R}\leq K\exp(L(t-\lfloor t\rfloor))/M$.
-The proof and details are left to Sec.
-C.2.
+The proof and details are left to Sec.˜C.2.
 Our sampler is in the same spirit as the Parareal simulation algorithms [54, 11, 10, 16]: it first obtains estimates at discrete steps and then runs fine-grained simulations for each interval.
@@
 Notably, the SDE simulation requires no additional NFEs and adds negligible computational overhead compared to neural-network inference, since it involves only simple arithmetic operations that can be executed efficiently on CPUs.
-For reference, we also present the standard EM solver in Sec.
-D.4.
-5 Experiments
-In this section, we extensively examine Neural MJD’s performance on synthetic and real-world time-series datasets, highlighting its applicability in business analytics and stock price prediction.
+For reference, we also present the standard EM solver in Sec.˜D.5.
 Table 1: Quantitative results on the synthetic dataset.
-Mean Winner-takes-all Probabilistic
+Mean Best-of-$K$ Probabilistic
 […]
@@
 Figure 3: Qualitative result on the synthetic dataset.
+5 Experiments
+In this section, we extensively examine Neural MJD’s performance on synthetic and real-world time-series datasets, highlighting its applicability in business analytics and stock price prediction.
 Baselines.
 We evaluate Neural MJD against a wide range of competitors, including statistical methods such as ARIMA [17], the BS model, and the MJD model.
-Additionally, we compare against learning-based approaches, including supervised learning models such as XGBoost [78], MLPs, and GCNs [79], as well as denoising diffusion models like DDPM [41], EDM [80], and flow matching (FM) [43].
-We include recent neural ODE or SDE based learning methods such as NJ-ODE [81], NeuralCDE [29] and LatentSDE [55] for comparisons.
+Additionally, we compare against learning-based approaches, including supervised learning models such as XGBoost [22], MLPs, GCNs [44], as well as denoising diffusion models like DDPM [36], EDM [39], and flow matching (FM) [55].
+We include recent neural ODE or SDE based learning methods such as NeuralCDE [41] and LatentSDE [51] for comparisons.
+NJ-ODE [35] was further included on the S&P 500 dataset.
 We also design a baseline model, Neural BS, which shares the same architecture as Neural MJD but omits the jump component.
@@
 Since some datasets contain graph-structured data as seen in the following section, we incorporate additional graph encoding steps based on Graphormer [87] to capture spatial features, which also justifies the inclusion of GCN as a baseline.
-Further details are provided in App.
-D.
+Further details are provided in App.˜D.
 Evaluation metrics.
 We employ Mean Absolute Error (MAE), Mean Squared Error (MSE), and the R-squared (R2) score as the primary evaluation metrics.
-To account for stochastic predictions, we run each stochastic models $10$ times and report results across three types of metrics: 1) Mean Metrics: used for deterministic models or to average the results of stochastic models; 2) Winner-Takes-All Metrics: we select the best prediction from ten stochastic samples to compute minMAE, minMSE, and maxR2; 3) Probabilistic Metrics: these metrics assess the likelihood of stochastic predictions and select the most probable outcome to calculate $p$-MAE, $p$-MSE, and $p$-R2.
+To account for stochastic predictions, we run each stochastic models $K=10$ times and report results across three types of metrics: 1) Mean Metrics: used for deterministic models or to average the results of stochastic models; 2) Best-of-$K$ (Oracle) Metrics: we select the best prediction from $K$ stochastic samples to compute minMAE, minMSE, and maxR2; 3) Probabilistic Metrics: these metrics assess the likelihood of stochastic predictions and select the most probable outcome to calculate $p$-MAE, $p$-MSE, and $p$-R2.
 We mark N/A for inapplicable metrics for certain methods.
-Please refer to App.
-D for more details.
+Please refer to App.˜D for more details.
 5.1 Synthetic Data
@@
 The data is split into 60% training, 20% validation, and 20% testing.
-Refer to App.
-D for details.
+Refer to App.˜D for details.
 Results.
-Tab.
-1 reports quantitative results on the jump-driven synthetic dataset.
+Tab.˜1 reports quantitative results on the jump-driven synthetic dataset.
 Learning-based methods outperform traditional statistical models (ARIMA, BS, and MJD), and our Neural MJD tops all three evaluation protocols, surpassing Neural BS thanks to its explicit jump modeling objective.
-Qualitatively, as shown in Fig.
-3, our Neural MJD generates larger, realistic jumps, while the baselines produce smoother but less accurate trajectories.
+Qualitatively, as shown in Fig.˜3, our Neural MJD generates larger, realistic jumps, while the baselines produce smoother but less accurate trajectories.
 Table 2: Quantitative results on the SafeGraph&Advan business analytics dataset.
-Metrics Mean Winner-takes-all Probabilistic
+Metrics Mean Best-of-$K$ Probabilistic
 […]
@@
 Neural MJD (ours) 54.1 4.18e04 0.549 42.3 3.19e04 0.565 53.0 4.10e04 0.550
+Table 3: Quantitative results on the S&P 500 stock dataset.
+Metrics Mean Best-of-$K$ Probabilistic
+[…]
+ARIMA 62.1 3.67e04 -0.863 N/A N/A
+BS 65.1 4.01e04 0.052 44.6 1.79e04 0.145 52.8 2.03e04 0.105
+MJD 64.3 3.58e04 0.092 40.7 1.22e04 0.235 49.7 1.67e04 0.112
+XGBoost 44.3 1.64e04 0.170 N/A N/A
+MLP 44.4 1.57e04 0.205 N/A N/A
+GCN 44.7 1.53e04 0.224 N/A N/A
+NJ-ODE 46.8 1.69e04 0.208 N/A N/A
+DDPM 42.2 1.88e04 0.235 36.8 8.42e03 0.470 N/A
+EDM 37.1 1.68e04 0.249 27.6 5.01e03 0.542 N/A
+FM 34.9 8.47e03 0.368 19.8 3.59e03 0.625 N/A
+NeuralCDE 42.8 1.46e04 0.201 N/A N/A
+LatentSDE 39.8 1.44e04 0.212 20.8 3.49e03 0.617 N/A
+Neural BS 31.6 4.32e03 0.781 12.6 8.04e02 0.959 22.3 2.19e03 0.889
+Neural MJD (ours) 15.4 1.36e03 0.953 4.3 1.46e02 0.995 13.6 1.08e03 0.963
 5.2 Real-World Business and Financial Data
@@
 2017) sets.
-Refer to App.
-D for further details about the datasets.
+Refer to App.˜D for further details about the datasets.
 Results.
-Tab.
-2 reports results on the SafeGraph&Advan dataset covering POIs revenue prediction, which is measured in dollars.
+Tab.˜2 reports results on the SafeGraph&Advan dataset covering POIs revenue prediction, which is measured in dollars.
 Denoising generative models (e.g., DDPM, EDM, FM) show strong performance, significantly outperforming simple supervised baselines like GCN.
-Neural MJD further improves upon the strong FM baseline, especially in winner-takes-all metrics, indicating better diversity and accuracy in generating plausible outcomes through simulated jumps.
+Neural MJD further improves upon the strong FM baseline, especially in Best-of-$K$ metrics, indicating better diversity and accuracy in generating plausible outcomes through simulated jumps.
 While denoising models support likelihood evaluation, their high computational cost—requiring hundreds of NFEs—makes them unsuitable for large datasets.
 In contrast, Neural MJD enables fast likelihood evaluation without such overhead, which enables the computation of probabilistic metrics.
-Tab.
-3 shows similar results on the S&P 500 dataset.
+Tab.˜3 shows similar results on the S&P 500 dataset.
 FM again outperforms conventional baselines, including ODE based NJ-ODE and NeuralCDE, and Neural MJD achieves the best overall performance, effectively capturing volatility and discontinuities in stock time-series data.
-Table 3: Quantitative results on the S&P 500 stock dataset.
-Metrics Mean Winner-takes-all Probabilistic
-[…]
-ARIMA 62.1 3.67e04 -0.863 N/A N/A
-BS 65.1 4.01e04 0.052 44.6 1.79e04 0.145 52.8 2.03e04 0.105
-MJD 64.3 3.58e04 0.092 40.7 1.22e04 0.235 49.7 1.67e04 0.112
-XGBoost 44.3 1.64e04 0.170 N/A N/A
-MLP 44.4 1.57e04 0.205 N/A N/A
-GCN 44.7 1.53e04 0.224 N/A N/A
-NJ-ODE 46.8 1.69e04 0.208 N/A N/A
-DDPM 42.2 1.88e04 0.235 36.8 8.42e03 0.470 N/A
-EDM 37.1 1.68e04 0.249 27.6 5.01e03 0.542 N/A
-FM 34.9 8.47e03 0.368 19.8 3.59e03 0.625 N/A
-NeuralCDE 42.8 1.46e04 0.201 N/A N/A
-LatentSDE 39.8 1.44e04 0.212 20.8 3.49e03 0.617 N/A
-Neural BS 31.6 4.32e03 0.781 12.6 8.04e02 0.959 22.3 2.19e03 0.889
-Neural MJD (ours) 16.8 1.22e03 0.938 5.2 1.61e02 0.992 15.0 1.02e03 0.948
-5.3 Ablation Study
+For completeness, we also report results for additional deterministic time-series baselines in Appendix D.3.
 Table 4: Ablation study on the effect of teacher forcing (TF) and Euler–Maruyama (EM).
-Mean Winner-takes-all Probabilistic
+Mean Best-of-$K$ Probabilistic
 […]
@@
 w. EM 85.6 0.397 79.4 0.423 84.4 0.405
-Table 5: Runtime comparison.
+Table 5: Runtime comparison of various methods.
 Model Train (ms) 1-run (ms) 10-run (ms)
@@
 Ours 183.5 166.8 179.2
-We perform ablation studies to evaluate (i) the training algorithm described in Alg.
-1 and (ii) the Euler-Maruyama with restart solver introduced in Sec.
-4.3.
+5.3 Ablation Study
+We perform ablation studies to evaluate (i) the training algorithm described in Alg.˜1 and (ii) the Euler-Maruyama with restart solver introduced in Sec.˜4.3.
 For the ablations, we use 10% of the SafeGraph&Advan business analytics training set for training and evaluate on the full validation set.
-The results are presented in Tab.
-5.
+The results are presented in Tab.˜5.
 Our training algorithm computes the MLE loss using the model predictions instead of ground truth, unlike teacher forcing.
@@
 Additionally, we empirically show that the vanilla EM solver results in higher variance and worse performance compared to our solver.
-Additionally, we compare the runtime of our method against various baselines in Tab.
-5.
+Additionally, we compare the runtime of our method against various baselines in Tab.˜5.
 Thanks to the efficient numerical simulation-based forecasting framework that does not increase NFEs, our models are particularly well-suited for efficient multi-run stochastic predictions.
@@
 Future work includes extending to more challenging data types like videos.
-7 Acknowledgements
+Acknowledgments and Disclosure of Funding
 This work was funded, in part, by the NSERC DG Grant (No.
@@
 Q.Y.
-is supported by UBC Four Year Doctoral Fellowships.
+is supported by UBC Four Year Doctoral Fellowship.
 Appendix
-[…]
-[…]
-l1
 Appendix A Derivations of the Stationary Merton Jump Diffusion Model
@@
 Our primary focus is on the case where the state variable $S$ is scalar, as is common in many studies.
-However, in Sec.
-4, we extend our analysis to the more general $\mathbb{R}^{d}$ setting.
+However, in Sec.˜4, we extend our analysis to the more general $\mathbb{R}^{d}$ setting.
 Notably, in our framework, we do not account for correlations among higher-dimensional variables.
@@
 Definition A.1.
-Lévy process [61, Definition 3.1] A càdlàg (right-continuous with left limits) stochastic process $(X_{t})_{t\geq 0}$ on $(\Omega,\mathcal{F},\mathbb{P})$ with values in $\mathbb{R}^{d}$ such that $X_{0}=0$ is called a Lévy process if it possesses the following properties:
+Lévy process [79, Definition 3.1] A càdlàg (right-continuous with left limits) stochastic process $(X_{t})_{t\geq 0}$ on $(\Omega,\mathcal{F},\mathbb{P})$ with values in $\mathbb{R}^{d}$ such that $X_{0}=0$ is called a Lévy process if it possesses the following properties:
   1. 1.
@@
 is a specific example of a Lévy process, as it comprises both a continuous diffusion component and a jump component.
-According to the Lévy–Itô decomposition [61, Proposition 3.7], any Lévy process can be expressed as the sum of a deterministic drift term, a Brownian motion component, and a pure jump process, which is represented as a stochastic integral with respect to a Poisson random measure.
+According to the Lévy–Itô decomposition [79, Proposition 3.7], any Lévy process can be expressed as the sum of a deterministic drift term, a Brownian motion component, and a pure jump process, which is represented as a stochastic integral with respect to a Poisson random measure.
 A.2 Explicit Solution to MJD
-To derive the solution to MJD in Eq. 2, based on [61, Proposition 8.14], we first apply Itô’s formula to the SDE:
+To derive the solution to MJD in Eq.˜2, based on [79, Proposition 8.14], we first apply Itô’s formula to the SDE:
 $\mathrm{d}f(S_{t},t)$ $=\frac{\partial f(S_{t},t)}{\partial t}\mathrm{d}t+b_{t}\frac{\partial f(S_{t},t)}{\partial S_{t}}\mathrm{d}t+\frac{\omega_{t}^{2}}{2}\frac{\partial^{2}f(S_{t},t)}{\partial S_{t}^{2}}\mathrm{d}t+\omega_{t}\frac{\partial f(S_{t},t)}{\partial S_{t}}\mathrm{d}W_{t}$
@@
 Here, $Y_{i}$ is the magnitude of the multiplicative jump.
-Therefore, integrating both sides of Eq.
-16, we get the final explicit solution for MJD model:
+Therefore, integrating both sides of Eq.˜16, we get the final explicit solution for MJD model:
 $\ln S_{t}-\ln S_{0}=(\mu-\lambda k-\frac{1}{2}\sigma^{2})t+\sigma W_{t}+\sum^{N_{t}}_{i=1}\ln Y_{i}.$ (17)
@@
 A.3 Likelihood Function of MJD
-For the log-likelihood derivation, given the conditional probability in Eq. 4, the log-likelihood of the MJD model can be expressed as:
+For the log-likelihood derivation, given the conditional probability in Eq.˜4, the log-likelihood of the MJD model can be expressed as:
 $\log P\left(\ln S_{t}|S_{0}\right)=$ $\log\sum_{n=0}^{\infty}P\left(N_{t}=n\right)P\left(\ln S_{t}|S_{0},N_{t}=n\right)$
@@
 $g(\omega)=i\omega k_{1}-\frac{\omega^{2}k_{2}}{2}+\frac{\omega^{3}k_{3}}{6}$... (24)
-According to [61, Proposition 3.13], the cumulants of the Lévy distribution increase linearly with $t$. Therefore, the first cumulant $k_{1}$ is the mean of the standard MJD:
+According to [79, Proposition 3.13], the cumulants of the Lévy distribution increase linearly with $t$. Therefore, the first cumulant $k_{1}$ is the mean of the standard MJD:
 $k_{1}(t)=\mathbb{E}\left[\ln\frac{S_{t}}{S_{0}}\right]=(\mu-\lambda k-\sigma^{2}/2+\lambda\nu)t$ (25)
@@
 $\text{Skewness}=k_{3}(t)=$ $\lambda(3\gamma^{2}\nu+\nu^{3})t$ (27)
-Excess $\text{Excess Kurtosis}=k_{4}(t)=$ $\lambda(3\gamma^{4}+6\nu^{2}\gamma^{2}+\nu^{4})t$ (28)
+$\text{Excess Kurtosis}=k_{4}(t)=$ $\lambda(3\gamma^{4}+6\nu^{2}\gamma^{2}+\nu^{4})t$ (28)
 Appendix B Derivations of the Non-stationary Merton Jump Diffusion Model
@@
 Definition B.1.
-Additive process [61, Definition 14.1] A stochastic process $(X_{t})_{t\geq 0}$ on $\mathbb{R}^{d}$ is called an additive process if it is càdlàg, satisfies $X_{0}=0$, and has the following properties:
+Additive process [79, Definition 14.1] A stochastic process $(X_{t})_{t\geq 0}$ on $\mathbb{R}^{d}$ is called an additive process if it is càdlàg, satisfies $X_{0}=0$, and has the following properties:
   1. 1.
@@
 the parameters governing drift, volatility, and jump intensity evolve over time, resulting in non-stationary increments.
-This violates the key stationarity property required for Lévy processes, as discussed in App.
-A.
+This violates the key stationarity property required for Lévy processes, as discussed in App.˜A.
 Consequently, the non-stationary MJD no longer falls within the Lévy process framework.
@@
 B.2 Explicit Solution to Non-stationary MJD
-To derive the explicit solution to the non-stationary MJD, according to [61, Proposition 8.19], we have the Itô formula for semi-martingales:
+To derive the explicit solution to the non-stationary MJD, according to [79, Proposition 8.19], we have the Itô formula for semi-martingales:
 $f(t,X_{t})-f(0,X_{0})$ $=\int^{t}_{0}\frac{\partial f(s,X_{s})}{\partial s}\mathrm{d}s+\int^{t}_{0}\frac{\partial f(s,X_{s-})}{\partial x}\mathrm{d}X_{s}+\frac{1}{2}\int^{t}_{0}\frac{\partial^{2}f(s,X_{s-})}{\partial x^{2}}\mathrm{d}[X,X]^{c}_{s}$
 $+\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[f(s,X_{s})-f(s,X_{s-})-\triangle X_{s}\frac{\partial f(s,X_{s-})}{\partial x}].$ (30)
-According to [61, Remark 8.3], for a function independent of time (i.e., $f(t,X_{t})=f(X_{t})$), when we have finite number of jumps, we can rewrite the above equation as:
+According to [79, Remark 8.3], for a function independent of time (i.e., $f(t,X_{t})=f(X_{t})$), when we have finite number of jumps, we can rewrite the above equation as:
 $f(X_{t})-f(X_{0})$ $=\int^{t}_{0}f^{\prime}(X_{s-})\mathrm{d}X_{s}^{c}+\frac{1}{2}\int^{t}_{0}f^{\prime\prime}(X_{s-})\mathrm{d}[X,X]^{c}_{s}+\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[f(X_{s})-f(X_{s-})],$
-where $X_{s}^{c}$ is the continuous part of $X_{s}$, and $[X,X]^{c}_{s}$ is the continuous quadratic variation of $X$ over the interval $[0,s]$.
+where $X_{s}^{c}$ is the continuous part of $X_{s}$, and $X,X]^{c}_{s}$ is the continuous quadratic variation of $X$ over the interval $[0,s]$.
 In our case, let $X_{t}=S_{t}$, and define $f(t,X_{t})=\ln S_{t}$, the corresponding derivatives are $\frac{\partial f(t,X_{t})}{\partial X_{t}}=\frac{\partial\ln S_{t}}{\partial S_{t}}=\frac{1}{S_{t}}$, and $\frac{\partial^{2}f(t,X_{t})}{\partial X_{t}^{2}}=\frac{1}{-S_{t}^{2}}$.
@@
 Thus, the jump contribution is $\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[f(X_{s})-f(X_{s-})]=\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[\ln(yS_{s-})-\ln(S_{s-})]=\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[\ln y]$.
-Since the jump process is driven by a Poisson random measure ${N}(\mathrm{d}t,\mathrm{d}y)$ on $[0,t]\times\mathbb{R}^{d}$, we can rewrite the sum over all jump times as an integral with respect to this measure.
+Since the jump process is driven by a Poisson random measure ${N}(\mathrm{d}t,\mathrm{d}y)$ on $0,t]\times\mathbb{R}^{d}$, we can rewrite the sum over all jump times as an integral with respect to this measure.
 When there are finitely many jumps on $[0,t]$, we have $\sum_{0\leq s\leq t,\triangle X_{s}\neq 0}[\ln y]=\int_{0}^{t}\int_{\mathbb{R}^{d}}\ln y{N}(\mathrm{d}s,\mathrm{d}y)$.
-Based on [61, Ch.
+Based on [79, Ch.
 14], even when the parameters (drift, volatility, jump intensity, etc.) are time-dependent, the non-stationary MJD remains a semi-martingale.
@@
 $\text{Var}\left[\ln(S_{t}/S_{0})\right]=\int_{0}^{t}\sigma_{s}^{2}\mathrm{d}s+\int_{0}^{t}\int_{\mathbb{R}^{d}}(\ln y)^{2}\lambda_{s}f_{Y}(s,y)\mathrm{d}y\mathrm{d}s.$ (35)
-Given the results for the general time-inhomogeneous system, one can directly substitute the coefficients into the discrete formulation implemented in Sec. 4.2 to obtain the corresponding results.
+Given the results for the general time-inhomogeneous system, one can directly substitute the coefficients into the discrete formulation implemented in Sec.˜4.2 to obtain the corresponding results.
 B.3 Likelihood Function of Non-stationary MJD
 Let $X_{t}=\ln S_{t}/\ln S_{0},{t\geq 0}$ be the log-return of the asset price $S_{t}$.
-Under the non-stationary MJD settings, $X_{t}$ is an additive process, therefore by the general property of additive process [61, Ch 14], the law of $X_{t}$ is infinitely divisible and its characteristic function is given by the Lévy–Khintchine formula:
+Under the non-stationary MJD settings, $X_{t}$ is an additive process, therefore by the general property of additive process [79, Ch 14], the law of $X_{t}$ is infinitely divisible and its characteristic function is given by the Lévy–Khintchine formula:
 $\mathbb{E}[\exp(iu\cdot X_{t})]=\exp\psi_{t}(u),$
@@
 When there is exactly one jump in $[0,t]$, the jump time $s_{1}$ is random.
-Given jump time $s_{1}$, in a time-inhomogeneous setting, the instantaneous probability of a jump at time $s_{1}$ is proportional to $\lambda_{s_{1}}$.
+Given jump time $s_{1}$, in a time‐inhomogeneous setting, the instantaneous probability of a jump at time $s_{1}$ is proportional to $\lambda_{s_{1}}$.
 According to the dynamics of non-stationary MJD, the continuous part of the log-return leads to a normal distribution with mean being $a_{1}=\ln S_{0}+\int^{s_{1}}_{0}\left(\mu_{s}-\lambda_{s}k_{s}-\frac{\sigma_{s}^{2}}{2}\right)\mathrm{d}s+\nu_{s_{1}}$, and variance being $b^{2}_{1}=\int_{0}^{s_{1}}\sigma_{s}^{2}\mathrm{d}s+\gamma^{2}_{s_{1}}$.
@@
 Appendix C Proofs of Theorem and Proposition
-C.1 Proof of Theorem 4.1
+C.1 Proof of Theorem˜4.1
 See 4.1 Before diving into the proof, we first introduce two important lemmas.
@@
 And $\Phi(x)$ is the cumulative distribution function (CDF) of the standard normal distribution and $\mathrm{sign}(\cdot)$ is the signum function.
-Lemma C.1 is particularly helpful in our proof below. We also acknowledge its foundation in an earlier work [87], which provides many insights and a profound amount of valuable knowledge on its own.
+Lemma˜C.1 is particularly helpful in our proof below. We also acknowledge its foundation in an earlier work [96], which provides many insights and a profound amount of valuable knowledge on its own.
 Lemma C.2 (Bounds on the Standard Normal CDF).
@@
 Proof.
-By the Mills’ ratio inequality for the Gaussian distribution [88], we have $01-\Phi(x)<\frac{\phi(x)}{x},\forall x>$ 0 .
-Using the identity $\Phi(-x)=1-\Phi(x)$ for $x>0$, we immediately obtain: $\Phi(-x)<\frac{\phi(x)}{x},\forall x>0.$ For $x<0$, substituting $-x$ into the previous bound and noting that $\phi(-x)=\phi(x)$, we obtain $\Phi(x)<\frac{\phi(x)}{|x|},\forall x<0.$
-Proof of Theorem 4.1.
+By the Mills’ ratio inequality for the Gaussian distribution [32], we have $1-\Phi(x)<\frac{\phi(x)}{x},\forall x>0.$ Using the identity $\Phi(-x)=1-\Phi(x)$ for $x>0$, we immediately obtain: $\Phi(-x)<\frac{\phi(x)}{x},\forall x>0.$ For $x<0$, substituting x-x into the previous bound and noting that $\phi(-x)=\phi(x)$, we obtain $\Phi(x)<\frac{\phi(x)}{|x|},\forall x<0.$
+Proof of Theorem˜4.1.
 Proof.
-The original likelihood objective in Sec. 4.2 is as follows:
+The original likelihood objective in Sec.˜4.2 is as follows:
 $P\left(\ln S_{t+\delta}|S_{t},{\mathcal{C}}\right)=\sum_{n=0}^{\infty}P\left(\Delta N=n\right)P\left(\ln S_{t+\delta}|S_{t},{\mathcal{C}},\Delta N=n\right)$ (42)
@@
 This completes the proof.
-C.2 Proof of Proposition 4.2
+C.2 Proof of Proposition˜4.2
 See 4.2
@@
 Step 2: Standard EM Results on Time-Inhomogeneous MJD SDEs
-Our paper considers time-inhomogeneous MJD SDEs, with parameters fixed within each interval $[\tau-1,\tau)$ ($\tau\in\mathbb{N}$, $\tau\geq 1$).
+Our paper considers time-inhomogeneous MJD SDEs, with parameters fixed within each interval $\tau-1,\tau)$ ($\tau\in\mathbb{N}$, $\tau\geq 1$).
 This happens to align with the Euler-Peano scheme for general time-inhomogeneous SDEs approximation.
@@
 We now discuss the error bound for the restarted EM solver, $\epsilon_{t}^{R}$.
-Thanks to explicit solutions for future states $\{S_{1},S_{2},\ldots,S_{T_{f}}\}$, we can analytically compute their mean $\mathbb{E}[S_{\tau}|\mathcal{C}]$, $\tau\geq 1$, based on Eq.
-12, which greatly simplifies the analysis.
-Using the restart mechanism in line 10 of Alg.
-2, we ensure that $\mathbb{E}[\bar{S}_{\tau}|\mathcal{C}]$ from our restarted EM solver closely approximates the true $\mathbb{E}[S_{\tau}|\mathcal{C}]$ at restarting times.
+Thanks to explicit solutions for future states $\{S_{1},S_{2},\ldots,S_{T_{f}}\}$, we can analytically compute their mean $\mathbb{E}[S_{\tau}|\mathcal{C}]$, $\tau\geq 1$, based on Eq.˜12, which greatly simplifies the analysis.
+Using the restart mechanism in line 10 of Alg.˜2, we ensure that $\mathbb{E}[\bar{S}_{\tau}|\mathcal{C}]$ from our restarted EM solver closely approximates the true $\mathbb{E}[S_{\tau}|\mathcal{C}]$ at restarting times.
 $\epsilon_{t}^{R}$ is significantly reduced when restart happens (when $t$ is an integer in our context for simplicity), then it grows again at the same rate as the standard EM method until the next restart timestep.
@@
 While both involve SDE-based representations of data, their problem formulations and learning objectives differ significantly.
-We illustrate the training loss computation pipeline for Neural MJD in Fig.
-4.
+We illustrate the training loss computation pipeline for Neural MJD in Fig.˜4.
 Notably, the loss computation can be processed in parallel across the future time-step horizon, eliminating the need for recursive steps during training.
@@
 For non-denoising models, normalization maps the data to the range $[0,1]$.
-In contrast, for denoising models (DDPM, EDM, FM), we scale the data to $[-1,1]$ to align with standard settings used in image generation.
+In contrast, for denoising models (DDPM, EDM, FM), we scale the data to [1,1][-1,1] to align with standard settings used in image generation.
 Importantly, normalization coefficients are derived solely from the training set statistics.
@@
 Additionally, ego graphs are constructed based on geodesic distances, where each POI serves as a central node connected to its 10 nearest neighbors.
-An visualization is shown in Fig.
-5.
+An visualization is shown in Fig.˜5.
 Specifically, we use POI area, brand name, city name, top and subcategories (based on commercial behavior), and parking lot availability as static features.
@@
 Based on the top category, we determine the maximum spending in the training data and use it to normalize the input data for both training and evaluation, ensuring a regularized numerical range.
-For training stability, we clip the minimum spending value to $0.01$ instead of 00 to enhance numerical stability for certain methods.
+For training stability, we clip the minimum spending value to $0.01$ instead of 0 to enhance numerical stability for certain methods.
 We adopt a sliding window approach with a stride of 1, using the past 14 days as input to predict spending for the next 7 days.
@@
 In total, it contains approximately 62K sequences for training, 5K for validation, and 15K for testing.
-To better distinguish the effects of different methods on the S&P 500 dataset, we use a adjusted R2 score $R^{2}=1-(1-R^{2}_{\text{reg}})\cdot\frac{n-1}{n-p-1}$, where $n$ is the sample size and we set the number of explanatory variables $p$ to be $(k-1)(n-1)/k$, where $k=70.0$.
-D.3 Limitations
+To better distinguish the effects of different methods on the S&P 500 dataset, we use a adjusted R2 score $R^{2}=1-(1-R^{2}_{\text{reg}})\cdot\frac{n-1}{n-p-1}$, where $n$ is the sample size and we set the number of explanatory variables $p$ to be (k1)(n1)/k(k-1)(n-1)/k, where $k=70.0$.
+D.3 Additional Deterministic Time-Series Baselines (Third-Party Implementations)
+For completeness, we also report results from third-party implementations of Autoformer [85], TiDE [27], and N-HiTS [20], provided by the NeuralForecast library (NIXTLA).
+Results in the table were produced with the publicly available NeuralForecast package on the same train/validation/test splits and identical data input (e.g.
+exogenous stock ticker information) as our main experiments, using the package’s default training settings without modification.
+Table 6: Quantitative results from NeuralForecast (NIXTLA) implementations on the S&P 500 stock dataset with tuned steps.
+Model MAE $\downarrow$ MSE $\downarrow$ $R^{2}$ $\uparrow$
+Autoformer 67.3 1.36e04 0.307
+TiDE 18.5 1.71e03 0.922
+N-BEATS 15.3 1.20e03 0.939
+N-HiTS 14.1 9.33e02 0.953
+TCN 14.4 9.73e02 0.951
+D.4 Limitations
 Our approach explicitly models discontinuities (jumps) in the time series.
@@
 Additionally, one potential extension is to design an adaptive mechanism that can automatically deactivate or regularize the jump component when the data do not exhibit significant jump behavior, thereby reducing unnecessary complexity and improving general performance on smooth series.
-D.4 Vanilla Euler Solver
+D.5 Vanilla Euler Solver
 Algorithm 3 Vanilla Euler-Maruyama Method
 0: Total solver steps $M$
-1:  ${\bm{C}}\sim\mathcal{D}_{\text{test}}$, with ${\bm{C}}=[S_{-T_{p}:0},C]$
-2:  $\{\mu_{\tau},\sigma_{\tau},\lambda_{\tau},\nu_{\tau},\gamma_{\tau}\}_{\tau=1}^{T_{f}}\leftarrow f_{\theta}({\bm{C}})$
+[…]
+[…]
 […]
@@
 10:return $\{\bar{S}_{t_{i}}\}_{i=1}^{M}$
-We present the standard Euler–Maruyama solver above, which is used in the ablation study for comparison with our restarted Euler solver.
+We present the standard Euler–Maruyama solver in Alg.˜3, which is used in the ablation study for comparison with our restarted Euler solver.
 Appendix E Impact Statement
@@
 We emphasize responsible deployment and continuous evaluation to mitigate inequalities and risks.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: The abstract and introduction accurately reflect our three main contribution, i.e., the time-inhomogeneous Neural MJD model, the ability of tractable learning, and empirical validation on synthetic and real-world datasets.
+The supporting theoretical derivations are provided in Sections 3,  4 and appendix (for long proofs), and the empirical results are provided in Section  5.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: The "Limitations" subsection is included in Appendix D.4.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: Every novel theorem provides its full set of assumptions, and complete proofs are provided in the appendix. The theorems drawn from standard textbooks are explicitly cited.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: We have clearly stated the training (Algorithm 1) and sampling (Algorithm 2) procedure in algorithm boxes, and the modeling and dataset details are described in both Section 5 and the appendix.
+Additionally, we will release our code once the paper is accepted.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24.
+Justification: We will release our code once the paper is accepted.
+For the dataset, the synthetic and S&P 500 dataset are publicly accessible, while the anonymized and aggregated SafeGraph&Advan data can be publicly purchased through Dewey platform.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29.
+Justification: We include full procedure for training (Algorithm 1) and sampling (Algorithm 2), outlining every step of our procedure.
+In Section 5 and Appendix D, we describe our dataset preprocessing in full—including the data-splitting strategy, normalization steps, and relevant hyperparameter settings—to ensure that our experimental setup can be replicated.
+Upon acceptance, we will release our source code and reproducible scripts.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34.
+Justification: Rather than conventional error bars, we report Best-of-$K$ and probabilistic metrics computed over multiple inference runs in Section 5, which capture the model’s predictive variability.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We include a detailed runtime comparison Tab. 5, and the specific GPU types used are listed in the Appendix D.1.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: Our work fully conforms with the NeurIPS ethics guidelines, using commercially licensed data responsibly, and ensuring no privacy, fairness, or other concerns arise from our work.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: The broader impacts are discussed as a separate section in the appendix.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54.
+Justification: Our work involves standard time-series forecasting on commercially licensed and public anonymized and aggregate datasets, and does not release any high-risk generative models or scraped data, so no special safeguards are required.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: Our models are clearly cited in both main text and the appendix, and dataset licenses are provided in the appendix.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64.
+Justification: We will release our code once the paper is accepted, which will be documented.
+For the dataset, the synthetic and S&P 500 dataset are publicly accessible, while the anonymized and aggregate SafeGraph&Advan data can be publicly purchased through Dewey platform (https://www.deweydata.io).
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: This project does not involve any crowdsourcing or human‐subject experiments.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: This project does not involve any human‐subject experiments.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: This project does not involve LLMs in any core method developments.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.